### v2

# DeepEdit: Knowledge Editing as Decoding with Constraints

Abstract
We propose a new perspective of knowledge editing (KE) for large language models (LLMs) that treats it as a constrained decoding problem. We design decoding constraints to regulate LLMs, ensuring coherence between reasoning steps when incorporating new knowledge. To enforce these constraints, we utilize a depth-first search to adaptively substitute new knowledge for the LLMs’ original reasoning steps, greedily seeking the optimal path of multi-hop reasoning with new knowledge. From this vantage, we propose DeepEdit: Depth-first Search-based Decoding for Knowledge Editing. DeepEdit improves the KE of LLMs by enhancing the conciseness, coherence, pertinence, and receptiveness of reasoning with new knowledge. DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions. In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which are designed to provide more precise and challenging assessments of KE approaches. Qualitatively, DeepEdit enables LLMs to produce more succinct reasoning outputs in accordance with new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
DeepEdit: Knowledge Editing as Decoding with Constraints
Yiwei Wang     Muhao Chen     Nanyun Peng     Kai-Wei Chang University of California, Los Angeles   University of California, Davis wangyw.evan@gmail.com https://wangywust.github.io/deepedit.io/
1 Introduction
Knowledge editing (KE) aims to enable large language models (LLMs) to produce outputs in accordance with the knowledge that is new or conflicted with the knowledge memorized by the LLMs. A challenging task to evaluate KE is the multi-hop question answering with new knowledge Zhong et al. (2023). In this task, LLMs should answer the questions corresponding to a chain of facts including new knowledge (see Figure 1(a)). It is essential for LLMs to deductively infer from the new knowledge in the multi-step reasoning in order to provide correct answers.
Figure 1: (a) An instance of multi-hop question answering with new knowledge. (b) The existing method MeLLo relies on the LLM generator to understand new knowledge by itself. (c) Our DeepEdit applies the constraints to the LLMs’ decoding so as to improve reasoning with new knowledge.
Figure 2: Our DeepEdit enforces LLMs to follow the decoding constraints of [Conciseness], [Coherence], [Receptiveness], and [Pertinence] so as to soundly incorporate new knowledge into LLMs’ reasoning.
Recent literature has mainly sought to perform KE through in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023a). However, in-context editing lacks direct control of the outputs; the new knowledge‘s influences are not traceable; and the performance suffers from the input-conflicting hallucinations Zhang et al. (2023b) (see Figure 1(b)).
Taking a closer look at the ground-truth output in Figure 1(a), we notice that: when reasoning with new knowledge, LLMs should orderly place new knowledge at the appropriate positions in the reasoning chain, i.e., the new knowledge should coherently bridge the adjacent reasoning steps, and avoid the memorized knowledge that is conflicted to new knowledge. This finding raises a natural question:
Can we directly control the decoding of LLMs’ to soundly incorporate the new knowledge?
The central idea of this paper is to view KE as a problem of constrained decoding. We design the decoding constraints, by following which LLMs can soundly incorporate new knowledge in LLMs reasoning. Specifically, we propose the following constraints: [Conciseness] prevents the redundant loops of reasoning; [Coherence] guarantees the coherence of adjacent steps; [Receptiveness] guarantees the LLMs’ awareness of new knowledge; [Pertinence] improves the relevance of reasoning to the input question (see Figure 4). These constraints improve LLMs’ reasoning with new knowledge from different perspectives as a whole.
Much research has proposed constrained decoding methods Meng et al. (2022c); Liu et al. (2023); Geng et al. (2023). However, classical constrained decoding methods usually require access to the token-wise distribution, which is not provided by many black-box LLMs like ChatGPT OpenAI (2022). In our work, to apply our constraints to the black-box LLMs’ decoding, we breakthrough this restriction by changing the granularity of decoding from tokens to the reasoning steps. We design all decoding actions over the step level, which bypasses the requirement of accessing the token-wise distributions.
At every iteration, we take one generated reasoning step and a few retrieved new facts as the step candidates. We verify the step candidates and filter out the ones that do not satisfy our constraints. Since there could be multiple candidates that satisfy all constraints every iteration. We operate the depth-first search that store all the valid step candidates in a stack and greedily search for the optimal path of LLMs’ multi-hop reasoning with new knowledge. We name our method as DeepEdit, the depth-first search-based decoding for knowledge editing.
The advantages of our DeepEdit on KE are three-folds. First, it does not need to retrain the LLMs like the classical model editing methods Zhu et al. (2020), which is complex and expensive. Second, the new knowledge directly contributes to the LLMs’ decoding by acting as the reasoning step candidates. This facilitates more effective reasoning with new knowledge. Third, the influence of new knowledge is monitored with decoding constraints, leading to more controllable and traceable reasoning with new knowledge. Overall, our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness.
DeepEdit can be flexibly applied to all the black-box LLMs. We evaluate DeepEdit on the KE benchmarks MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al. (2023); Meng et al. (2022a). The latter two are new benchmarks that we build to provide more precise and challenging assessments of KE methods. Qualitatively, DeepEdit leads to more succinct and faithful reasoning of LLMs that soundly incorporate new knoweldge, as shown in Figure 8. We also observe substantial quantitative improvements in the KE question-answering accuracy achieved by DeepEdit. Overall, DeepEdit improves the KE for the strong LLMs text-davinci-003, gpt-3.5-turbo-instruct, and llama2 by a significant margin OpenAI (2022); Touvron et al. (2023).
2 Related Work
Constrained Decoding
Constrained decoding Qin et al. (2020, 2022); Lu et al. (2020, 2021); Kumar et al. (2022); Liu et al. (2023); Geng et al. (2023) is a classical natural language processing topic, which is mainly applied to control the lexicon, sentiment or style of the generated text. FUDGE Yang and Klein (2021) employs weighted decoding (WD) by directly adjusting the probabilities of the vocabulary distribution with an auxiliary classifier probability. NADO Meng et al. (2022c) suggests sampling from a vocabulary distribution similar to an approximation of the sequence-level oracle. GeDi Krause et al. (2020) and DExperts Liu et al. (2021) also adopt the weighted decoding approach through the training of generative classifiers. Constrained decoding methods generally require access to the LLMs’ token-wise distribution. Different from them, our DeepEdit avoids the requirements of distribution access by decoding at the reasoning step level. We extend the applications of constrained decoding to KE. We design the semantic constraints to soundly incorporate new knowledge in LLMs’ reasoning, and verify the complex semantic constraints with LLMs.
Figure 3: Comparison of different KE methods. (upper) Model editing was the most popular KE paradigm, which edits the model parameters to update knowledge. (middle) Recently, in-context editing has become popular, which relies on the LLMs’ capability of utilizing new knowledge in the prompts. (lower) We propose DeepEdit for KE, which directly controls the LLMs’ decoding to follow new knowledge.
Knowledge Editing
Previous studies have explored multiple methods for editing the knowledge of LLMs by introducing new knowledge into static model artifacts Zhu et al. (2020); Sotoudeh and Thakur (2019); Dai et al. (2021); Hase et al. (2021); Zhou et al. (2023); Dong et al. (2022); Huang et al. (2023). Some of these approaches involve identifying and modifying model weights associated with specific concepts Meng et al. (2022a, b); Dai et al. (2021), as well as rapid adaptation facilitated by a compact auxiliary editing network Mitchell et al. (2022a); De Cao et al. (2021); Mitchell et al. (2021). Given the fast-growing parameter sizes of large language models (LLMs) Zhao et al. (2023), frequently updating LLMs with new knowledge through retraining is more and more expensive. Hence, it is vital to effectively edit the LLMs’ knowledge without retraining.
Recent work observes the superior KE performance of in-context editing. MeLLo Zhong et al. (2023) designs a single prompt to alternately conduct the text generation and new knowledge analysis using the LLM generator. Cohen et al. (2023) propose to append the new knowledge the head of the input prompt. In-context editing lacks direct control of the LLMs’ outputs and how the new knowledge influences the outputs is not traceable. Different from them, we propose the new decoding method DeepEdit for KE, which effectively augments LLMs on KE during the decoding stage. We illustrate a comparison between DeepEdit and existing KE methods in Figure 3.
Specifically, in our work, we consider a challenging setting of knowledge editing: the multi-hop question answering (QA) with knowledge editing Zhong et al. (2023). In-context learning methods combined with an optional retrieval module have become a popular approach in handling multi-hop QA tasks Press et al. (2022); Yao et al. (2022); Khattab et al. (2022); Mitchell et al. (2022b). We propose four types of decoding constraints to enhance question answering with KE. Our work improves multi-hop QA with KE by applying constraints to the reasoning of LLMs, which directly controls LLMs’ outputs to follow new knowledge.
Figure 4: Comparison between classical constrained decoding (at the token level) (upper) and our decoding at the reasoning step level (lower). Differences: 1) The token candidates is the predefined vocabulary, while the step candidates include a generated step and the retrieved facts. 2) The token candidates are ranked by generators’ outputs, while the step candidates are filtered by the constraints.
Self-refinement of Large Language Models
Some work has explored to use LLMs to revise the output of themselves in dialogue response, sentiment reversal, and mathematical reasoning Madaan et al. (2023); Shridhar et al. (2023b, a); Welleck et al. (2022); Zheng et al. (2023b); Madaan et al. (2023). However, it is difficult to rely on LLMs to accommodate the new knowledge by themselves in the self-refinement, since the new knowledge is generally conflicted with the parametric knowledge memorized by LLMs. Different from them, we do not rely on LLMs to revise the outputs by themselves. We directly control the decoding to soundly incorporate new knowledge with decoding constraints. We modify the LLMs’ outputs using knowledge retrieval and depth-first search to satisfy the constraints. We hope that our work demonstrates more potential to conduct KE for LLMs in the decoding stage.
3 Methodology
Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) Zhong et al. (2023), where LLMs should answer the questions corresponding to a chain of facts including new knowledge. This task assumes a database storing all new facts that LLMs should follow, and only a few facts can influence the ground-truth answer to a specific question.
Our KE idea is to first design the decoding constraints that can regulate LLMs to soundly incorporate new knowledge in reasoning and then enforce LLMs’ decoding to follow the constraints. In this section, we will introduce how we incorporate our idea into the design of a new KE method DeepEdit (Depth-first Search based Decoding for Knowledge Editing) to improve the KE of black-box LLMs.
3.1 Decoding at the Reasoning Step Level
Many classical constrained decoding methods Zhang et al. (2023a); Meng et al. (2022c) use an auxiliary model to adjust the token-wise distribution during decoding. They manipulate the distribution before every token’s generation to satisfy the constraints. Therefore, they require accessibility to the LLMs’ output distribution, which is not feasible for many black-box LLMs.
In this work, we change the granularity of decoding from the token level to the reasoning step level. Our decoding is designed at the step level: instead of taking the tokens predefined in a fixed vocabulary as the candidates, we take the reasoning steps as the candidates. At every iteration, we let LLMs to produce one reasoning step and retrieve a fixed number of new facts in parallel to the reasoning step as the step candidates. The new fact retrieval is based on the nearest neighbor search with the sentence-level embeddings given by BERT. The motivation of this design is that the new facts that are semantically close to the generated facts are more likely to satisfy the constraints.
Taking the single generated step and the new facts as the step candidates, we filter a valid candidate and take it as the new reasoning step. Note that there may exist more than one candidates that satisfy the constraints. We will introduce how we design the decoding constraints for KE in Section 3.2, how we verify the constraints in Section 3.3 and Section 3.4, and how we process the multiple valid step candidates to search for the optimal path of LLMs’ multi-hop reasoning in Section 3.5.
Our decoding design bypasses the requirement of accessing the token-wise distributions, since all the decoding actions are defined over the step level. Our decoding strategy is free to be applied to black-box LLMs in order to produce satisfactory outputs in accordance with new knowledge. The details are introduced next.
Figure 5: (left) At every decoding iteration, we store the reasoning step candidates (including the generated step and the retrieved knowledge) that satisfy the local constraints in a stack. (right) The global verifier determines the feasibility of the whole reasoning process by checking whether the answer is found. If not feasible, we pop the status stack to use another reasoning subprocess that satisfies the local verifier.
3.2 Designing Decoding Constraints for Knowledge Editing
To correctly answer multi-hop questions in accordance with new knowledge Zhong et al. (2023), LLMs need to generate a sequence of reasoning steps to find the answer. We propose the following constraints for the LLMs’ reasoning with KE.
  1. 1.
    [Conciseness] Every reasoning step should be different from the prior ones.
  2. 2.
    [Coherence] Every reasoning step should be relevant to the last reasoning step.
  3. 3.
    [Receptiveness] Every reasoning step cannot contradict to any new knowledge.
  4. 4.
    [Pertinence] Every reasoning step should be relevant and helpful to finding the answer.
The above constraints are not automatically satisfied by LLMs when reasoning with new knowledge. However, they are essential to helping LLMs to find the correct answers in accordance with new knowledge. Therefore, in our work, we propose a new decoding method to enforce the LLMs to follow the above constraints so as to improve the KE performance. We first analyze the constraints in detail.
Among the above four constraints, [Conciseness] guarantees that the LLMs’ reasoning is not in a loop. [Coherence] guarantees the coherence of reasoning, while [Pertinence] guarantees the relevance of every reasoning step to the question. [Receptiveness] ensures that LLMs can utilize new knowledge for reasoning. We visualize different constraints in Figure 4. Among the constraints, we regard [Conciseness], [Coherence], and [Receptiveness] as local ones: we can verify whether they are satisfied given only partial reasoning steps. The [Pertinence] constraint is regarded as a global constraint: we can verify it based on a complete reasoning process.
By following the above constraints, LLMs can orderly place new knowledge at appropriate positions in the reasoning chain and avoid the memorized knowledge that is conflicted with new knoweledge. Next, we will introduce how we verify the above constraints and control the LLMs’ decoding to satisfy the constraints.
3.3 Verifier for Local Constraints
To verify the [Conciseness] constraint, by default, whenever a new reasoning step is given, we may compare it with all the prior reasoning steps. However, this is no doubt an expensive solution at the cost of $\mathcal{O}(r^{2})$ times of verification given as many as $r$ reasoning steps.
We observe that the repeated reasoning step generally occurs in the adjacent steps. Therefore, to reduce the computational costs, we propose an approximate verification method to verify the constraint of [Conciseness]. Specifically, for every reasoning step, we only compare it with the last reasoning step, so that the whole time of doing the verification is $\mathcal{O}(r)$.
Figure 6: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k. The chosen step candidate in every iteration that satisfies all the constraints are in green. The dropped step candidates are followed by the constraints that they do not satisfy.
To satisfy the [Coherence] constraint, we need to evaluate whether the current reasoning step is relevant to the last reasoning step. However, it is challenging to directly and precisely evaluate the relevance through LLMs. We use the LLM based verifiers equipped with a hand-crafted prompt to verify whether the adjacent reasoning steps are semantically equivalent or related to verify the constraints of [Conciseness] and [Coherence] respectively.
Verifying the [Receptiveness] constraint, by default, requires determining whether the contradiction exists in the pair of the reasoning step and all the items in the basis of new knowledge. The time complexity is $\mathcal{O}(rk)$, where $k$ is the number of new knowledge items and a large value in practice. Therefore, to reduce the time complexity, we retrieve the relevant knowledge items that are semantically closest to the reasoning step and then assess the contradictions. The idea is that if two sentences are contradictory to each other, they must first be semantically relevant. In this way, we reduce the time complexity from $\mathcal{O}(rk)$ to $\mathcal{O}(r\log(k))$. Also, we design a LLM-based verifier with a hand-crafted prompt to assess whether the retrieved facts are contradictory to the generated reasoning step.
3.4 Verifier for Global Constraints
Whether a reasoning step is helpful in finding the final answer is hard to justify until the final answer is found. Therefore, we only verify the [Pertinence] constraint globally for each reasoning step given the full completed reasoning process. When the intermediate reasoning steps are not helpful but even barrier the final answer, the reasoning likely takes more steps to find the final answer. Based on this prior knowledge, we set the maximum reasoning steps as $r$, so that the reasoning is terminated when an answer is not found given more than $r$ steps. We set $r=5$ by default in our experiments.
3.5 Decoding with Depth-first Search
As we have introduced in Section 3.1, there may exist multiple step candidates that satisfy all the constraints. In this case, how can we decide which step is the next step?
To answer this question, let’s first recall how LLMs how to decide which is next token. The most simple method is the greedy decoding, which just takes the highest logit token candiate as the next token. In our case, since we do not have scores to rank step candidates but only the discrete constraint verifications to filter the candidates, we cannot simply take the “best” step as the next reasoning step. Can we have other alternate solutions?
Let’s see a more advanced token-level decoding method: beam search. Beam search takes multiple candidate tokens as the next tokens and store them in a stack to do a best-first search in the token sequence generation. We borrow the idea of the beam search to process the multiple valid step candidates with the depth-first search. Specifically, we store all the step candidates that satisfy the constraints in a stack, as shown in Figure 5. Then, in every iteration, we pop the reasoning step stored at the head of the stack as the next reasoning step and do further reasoning based on it. We name our method as DeepEdit: depth-first search based decoding for knowledge editing. We visualize our DeepEdit in Figure 1(c). Also, we present the iteration-wise outputs of our DeepEdit on an instance in Figure 6.
The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al. (2023). Different from them, our DeepEdit’s depth-first search is to enforce the LLMs’ decoding to satisfy the constraints derived for KE from a global perspective.
4 New Benchmarks for More Precise and Challenging Evaluation
We provide two new benchmarks for the evaluation of KE methods: MQuAKE-2002 and MQuAKE-hard. Both are built from the MQuAKE Zhong et al. (2023), a recent KE benchmark that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge. Every instance in MQuAKE includes a multi-hop question and one to four related edited facts, every of which can alter the ground-truth answer to the question. Zhong et al. (2023) suggests using a randomly sampled 3,000 instance subset of MQuAKE to do the evaluation, which reduces the experimental costs. This evaluation benchmark is termed as MQuAKE-3k.
4.1 Annotation Mistakes in One-Third Instances of MQuAKE-3k
There are two issues of using MQuAKE-3k to evaluate KE. The first issue is that the new knowledge from different instances MQuAKE-3k can cause conflicts and mistakes to the ground-truth answers. In other words, the ground-truth answer from instance A can be altered by the new knowledge from another instance B. We show an example of knowledge conflicts in MQuAKE-3k in Figure 7. These knowledge conflicts will make the ground-truth answers incorrect given the new knowledge that are conflicted with the answers because the inference on every instance would retrieve the new knowledge from all instances. We observe that 998 instances’ ground-truth answers are broken by the new knowledge from other instances.
4.2 New Benchmark MQuAKE-2002 for More Precise Evaluation
To address the issue of annotation mistakes in MQuAKE-3k, we provide a new subset of MQuAKE-3k, which does not have any knowledge conflict across instances. This subset includes 2,002 instances, so we term it as MQuAKE-2002. We filter out the instances of which the ground-truth answers are broken by the new knowledge from other instances to produce MQuAKE-2002. Compared with MQuAKE-3k, our MQuAKE-2002 provides a more precise evaluation for knowledge editing methods, since it remove the annotation mistakes due to knowledge conflicts across instances. The data statistics of MQuAKE-2002 is provided in Table 1.
4.3 New Benchmark MQuAKE-hard for More Challenging Evaluation
The second issue of MQuAKE-3k is that more than 66% instances in MQuAKE-3k only contain at most two edited facts that influence the answers, which are not challenging enough to evaluate the knowledge editing methods on handling multiple edited facts that can alter the ground-truth reasons. we construct a more challenging subset of MQuAKE by selecting the instances that contain the highest number of edited facts per instance. We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts. MQuAKE-hard has no overlap with MQuAKE-3k. We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts. MQuAKE-hard has no overlap with MQuAKE-3k. The data statistics of MQuAKE-hard is also provided in Table 1.
Benchmark # Instances # Hops per Instance # Edited Facts per Instance # Conflicted Instances MQuAKE-3k Zhong et al. (2023) 3,000 3.0 2.0 998 MQuAKE-2002 2,002 2.7 2.2 0 MQuAKE-hard 4,29 4.0 4.0 0
Table 1: Data statistics of different benchmarks. # Conflicted Instances represent the number of instances of which the ground-truth labels are affected by the new knowledge from other instances. An example is shown in Figure 7.
Figure 7: An example of knowledge conflicts across instances in MQuAKE-3k. New knowledge (in red) in instance 382 is conflicted with the ground-truth reasoning step (in green) of instance 22.
5 Experiments
In this section, we evaluate the KE performance of our DeepEdit method when applied to LLMs, and compare against a strong baseline MeLLo Zhong et al. (2023). Our experimental settings closely follow those of the previous work Zhong et al. (2023) to ensure a fair comparison.
5.1 Experimental Settings
We take the following KE benchmarks MQuAKE-3k Zhong et al. (2023), MQuAKE-2002, and MQuAKE-hard for evaluation. The introduction of the above benchmarks are provided in Section 4. We follow the existing work Zhong et al. (2023) to use accuracy as the main evaluation metric. For every instance, we test the accuracy of the first question in the instance, since different questions in the same instance are semantically equivalent and correspond to the same ground-truth answer. We follow Zhong et al. (2023) to set the KE batch size as 100 and full batch for evaluation. The KE batch size means the number of instances that provide new knowledge for retrieval.
Compared methods. We take the strong KE methods into comparison: MEND Mitchell et al. (2021), MEMIT Meng et al. (2022b), IKE Zheng et al. (2023a), ICE Cohen et al. (2023), and MeLLo Zhong et al. (2023). MEND produces weight updates by transforming the raw fine-tuning gradients given an edited fact. MEMIT updates feedforward networks in a range of layers to encode the edited facts. IKE and ICE conduct KE by adding new knowledge to the input prompts. MeLLo is the state-of-the-art KE method that prompts the LLMs to generate subquestions and append new knowledge to the prompts. The former two KE methods are only applicable to white-box LLMs, while the latter three can fit the block-box LLMs.
Method MQuAKE-3k MQuAKE-hard batch size of 100 full batch batch size of 100 full batch llama2-7b w/ MEND Mitchell et al. (2021) 7.2 3.9 0.5 0.2 llama2-7b w/ MEMIT Meng et al. (2022b) 7.9 4.2 0.9 0.5 llama2-7b w/ IKE Zheng et al. (2023a) 12.5 6.2 1.2 0.5 llama2-7b w/ ICE Cohen et al. (2023) 12.9 6.3 1.4 0.7 llama2-7b w/ MeLLo Zhong et al. (2023) 16.3 10.8 1.9 1.6 llama2-7b w/ DeepEdit (Ours) 27.6 11.2 14.5 7.0 gpt-3.5 w/ IKE Zheng et al. (2023a) 15.5 8.3 1.6 1.2 gpt-3.5 w/ ICE Cohen et al. (2023) 16.1 8.5 2.1 1.4 gpt-3.5 w/ MeLLo Zhong et al. (2023) 29.9 20.0 3.7 1.6 gpt-3.5 w/ DeepEdit (Ours) 47.2 38.0 51.3 48.0 text-davinci-003 w/ IKE Zheng et al. (2023a) 18.2 8.9 1.9 1.2 text-davinci-003 w/ ICE Cohen et al. (2023) 19.0 8.6 2.3 1.2 text-davinci-003 w/ MeLLo Zhong et al. (2023) 32.2 25.6 4.7 2.3 text-davinci-003 w/ DeepEdit (Ours) 49.1 38.8 52.0 49.2
Table 2: Experimental results (accuracy; %) on the dataset MQuAKE-3k and MQuAKE-hard. The batch size means the number of instances providing the edited facts for knowledge retrieval. gpt-3.5 stands for the model gpt-3.5-turbo-instruct. The best KE result on every LLM is highlighted in bold font.
Method MQuAKE-2002 gpt-3.5 w/ MeLLo 25.1 gpt-3.5 w/ DeepEdit 53.7
Table 3: Experimental results (accuracy; %) of the full-batch knowledge editing on the benchmark MQuAKE-2002.
Verifier [Conciseness] [Coherence] [Receptiveness] llama2-7b 82.3 77.9 86.3 gpt-3.5 92.8 91.5 92.2
Table 4: Experimental results (verification accuracy; %) of our verifiers for different constraints on the dataset MQuAKE-3k.
gpt-3.5 w/ DeepEdit 49.1 $\Delta$ w/o [Conciseness] 46.1 $\downarrow 6\%$ w/o [Coherence] 42.9 $\downarrow 13\%$ w/o [Receptiveness] 45.2 $\downarrow 8\%$ w/o [Pertinence] 46.6 $\downarrow 5\%$
Table 5: Experimental results (accuracy; %) on the dataset MQuAKE-3k of DeepEdit with KE batch size of 100.
Model configuration. We set the hyper-parameters of the baseline methods as suggested by their papers. We apply the KE methods to the strong LLM text-davinci-003, gpt-3.5-turbo-instruct (short as gpt-3.5), and llama2-7b (short as llama2) OpenAI (2022); Touvron et al. (2023) for evaluation. The former two are black-box LLMs and the latter one is a white-box LLM. By default, we set the decoding temperature as 0.0 to minimize the randomness of LLMs’ outputs. The DeepEdit’s verifiers use the same LLM as the generator.
Figure 8: A case study for MeLLo and our DeepEdit on the dataset MQuAKE-3k. The new knowledge are in green. We report the prediction of text-davinci-003 with MeLLo and our DeepEdit.
5.2 Overall Performance
We incorporate DeepEdit with several choices of LLMs to test its effectiveness on editing LLMs’ knowledge. We report the test accuracy on MQuAKE-3k and MQuAKE-hard in Table 2. When the KE batch size is 100, our DeepEdit improves the accuracy of question answering with KE by 69%, 58%, and 52% over MeLLo on the LLMs llama2, gpt-3.5-turbo-instruct, and text-davinci-003 respectively over the strong baseline method MeLLo on MQuAKE-3k. On MQuAKE-hard, our DeepEdit outperforms the strong baseline method MeLLo by $\times$7, $\times$12, and $\times$10 times in terms of question answering accuracy with KE on the LLMs llama2, gpt-3.5, and text-davinci-003 respectively.
The experimental results on MQuAKE-2002 is shown in Table 3. When the KE batch size is full batch, our DeepEdit still outperforms the baseline methods and improves the KE accuracy by a significant margin. Our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness. The above advantages are especially important when the LLMs have to utilize multiple new knowledge to get the correct answers, since more new knowledge requires the LLMs to be correctly aware of the new knowledge on more reasoning steps. The correct knowledge incorporation with our decoding constraints significantly improves the KE performance given more new knowledge. Overall, our DeepEdit achieves substantial improvements on KE for LLMs and consistently outperforms the baseline methods on different benchmarks.
5.3 Analysis on Verifiers
We conduct deep analysis on the different verifiers of DeepEdit. We first test the accuracy of different verifiers to verify the corresponding constraint. We randomly sample 4,389 pairs of reasoning steps from the ground-truth reasoning steps of MQuAKE as the positive pairs that satisfy the constraints and randomly pairing the same number of steps from different questions as the negative examples. We evaluate different verifiers based on llama2 and gpt-3.5 on the sampled reasoning step pairs. We report the verification accuracy in Table 4. We observe high verification accuracy of around 80% and over 90% on different constraints achieved by llama-2 and gpt-3.5 respectively. Also, we observe that the performance of verifiers based on llama2 and gpt-3.5 are positively related to the LLMs’ performance with our DeepEdit, since the effective constraint assessment forms the basis of effective reasoning of LLMs with our DeepEdit.
5.4 Ablation Study
We conduct ablation studies on DeepEdit to empirically examine the contribution of its main technical components. We evaluate the influence of different constraints on DeepEdit. We report the experimental results of the ablation study of constraints in Table 5. We observe that removing each constraint leads to substantial performance degradation in different degrees. This indicates that different constraints contribute to more effective KE from the perspectives of progress, relevance, coherence, and awareness as a whole, and removing any of them will lead to more mistakes in the reasoning of LLMs. The above results also validate that the LLMs’ outputs cannot automatically satisfy the constraints that represent the desired properties without our DeepEdit.
5.5 Case Study
Figure 8 gives a qualitative comparison example between MeLLo and our DeepEdit on MQuaKE-3k. The results show that the strong KE method MeLLo makes mistakes on KE. For example, given the question “The writer of the novel “1984” died in which continent? ”, MeLLo is unable to generate the helpful sub-question to find the final correct answer. In contrast, our DeepEdit, based on the constraints of [Conciseness], [Coherence], and [Receptiveness], correctly find the relevant new knowledge “Bucharest is located in the continent of North America.” and obtain the correct answer North America with respect to the new knowledge.
Taking a closer look, we observe that our DeepEdit’s induced reasoning is much more succinct than MeLLo, since we effectively filter out the irrelevant and redundant information during decoding. Because longer and noisy textual information can lead to more serious hallucinations and higher difficulty of reasoning, our DeepEdit exhibits a stronger capability of KE for LLMs. Overall, our DeepEdit method enforces LLMs’ outputs to follow our decoding constraints so as to soundly incorporate new knowledge in reasoning, which leads to the correct predictions in Figure 8.
6 Conclusion
We explore a new paradigm for knowledge editing of black-box LLMs: decoding with constraints. We have designed a new decoding method DeepEdit to improve the knowledge editing of LLMs. We enforce the LLM’s outputs to satisfy the desired properties with decoding at the reasoning step level. Based on the DeepEdit, we can improve the LLMs’ reasoning with new knowledge though higher conciseness, coherence, pertinence, and receptiveness. Our DeepEdit controls LLMs’ reasoning to soundly incorporate new knowledge by satisfying our constraints. In addition to DeepEdit, we provide two new benchmarks: MQuAKE-2002 and MQuAKE-hard to provide more precise and challenging evaluation for KE methods respectively. Extensive experiments demonstrate the effectiveness of our DeepEdit. Future work includes analyzing the effects of other constraints that can improve knowledge editing for large language models.

### v3

@@
 Abstract
-We propose a new perspective of knowledge editing (KE) for large language models (LLMs) that treats it as a constrained decoding problem.
-We design decoding constraints to regulate LLMs, ensuring coherence between reasoning steps when incorporating new knowledge.
-To enforce these constraints, we utilize a depth-first search to adaptively substitute new knowledge for the LLMs’ original reasoning steps, greedily seeking the optimal path of multi-hop reasoning with new knowledge.
-From this vantage, we propose DeepEdit: Depth-first Search-based Decoding for Knowledge Editing.
-DeepEdit improves the KE of LLMs by enhancing the conciseness, coherence, pertinence, and receptiveness of reasoning with new knowledge.
-DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions.
-In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which are designed to provide more precise and challenging assessments of KE approaches.
-Qualitatively, DeepEdit enables LLMs to produce more succinct reasoning outputs in accordance with new knowledge.
+Answering multi-hop questions involving new knowledge is a challenging task in evaluating large language models’ (LLMs) knowledge editing (KE) methods.
+This task is rather difficult because the LLMs’ hallucinations on new knowledge would harm the logical coherence of LLMs’ multi-hop reasoning and lead to incorrect answers.
+To address this issue, we design decoding constraints to “regulate” LLMs’ reasoning, enhancing logical coherence when incorporating new knowledge.
+We incorporate the constraints into a new KE framework: DeepEdit ( Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs to generate coherent reasoning chains with new knowledge through a depth-first search.
+Our search selects the most important knowledge that satisfies our constraints as the reasoning step to efficiently increase the reasoning depth.
+In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which provide more precise and challenging assessments of KE approaches.
+Qualitatively, DeepEdit enables LLMs to produce succinct and coherent reasoning chains involving new knowledge.
 Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@
 In this task, LLMs should answer the questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
-It is essential for LLMs to deductively infer from the new knowledge in the multi-step reasoning in order to provide correct answers.
 Figure 1: (a) An instance of multi-hop question answering with new knowledge.
-(b) The existing method MeLLo relies on the LLM generator to understand new knowledge by itself.
-(c) Our DeepEdit applies the constraints to the LLMs’ decoding so as to improve reasoning with new knowledge.
-Figure 2: Our DeepEdit enforces LLMs to follow the decoding constraints of [Conciseness], [Coherence], [Receptiveness], and [Pertinence] so as to soundly incorporate new knowledge into LLMs’ reasoning.
+(b) Prior KE methods, e.g., MeLLo Zhong et al.
+(2023), suffers from the hallucinations of LLM generators.
+(c) Our DeepEdit controls LLMs with decoding constraints to produce coherent reasoning chains involving new knowledge.
+Figure 2: Our DeepEdit controls LLMs’ reasoning to follow the constraints of Conciseness, Coherence, Receptiveness, and Pertinence so as to soundly incorporate new knowledge into LLMs’ question answering.
 Recent literature has mainly sought to perform KE through in-context editing Zhong et al.
@@
 (2023); Zheng et al.
-(2023a).
-However, in-context editing lacks direct control of the outputs; the new knowledge‘s influences are not traceable; and the performance suffers from the input-conflicting hallucinations Zhang et al.
-(2023b) (see Figure 1(b)).
-Taking a closer look at the ground-truth output in Figure 1(a), we notice that: when reasoning with new knowledge, LLMs should orderly place new knowledge at the appropriate positions in the reasoning chain, i.e., the new knowledge should coherently bridge the adjacent reasoning steps, and avoid the memorized knowledge that is conflicted to new knowledge.
-This finding raises a natural question:
+(2023), which puts new knowledge in the input context to fulfill KE.
+However, the input-conflicting hallucinations on the in-context new knowledge would harm the logical coherence of multi-step reasoning and lead to incorrect answersZhang et al.
+(2023).
+Taking a closer look at the ground-truth output in Figure 1(a), we notice that: the adjacent reasoning steps should be logically coherent, and new knowledge should participate in the reasoning chain at appropriate positions.
+This raises a natural question:
 Can we directly control the decoding of LLMs’ to soundly incorporate the new knowledge?
@@
 We design the decoding constraints, by following which LLMs can soundly incorporate new knowledge in LLMs reasoning.
-Specifically, we propose the following constraints: [Conciseness] prevents the redundant loops of reasoning; [Coherence] guarantees the coherence of adjacent steps; [Receptiveness] guarantees the LLMs’ awareness of new knowledge; [Pertinence] improves the relevance of reasoning to the input question (see Figure 4).
-These constraints improve LLMs’ reasoning with new knowledge from different perspectives as a whole.
-Much research has proposed constrained decoding methods Meng et al.
-(2022c); Liu et al.
-(2023); Geng et al.
+Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted memorized knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 2).
+A multi-step reasoning chain that leads to the correct answer should satisfy the above constraints.
+We incorporate the above constraints into a new KE framework: DeepEdit ( Depth-first Search-based Constrained Decoding for Knowledge Editing).
+When producing every reasoning step, DeepEdit searches over the parametric and new knowledge to find the ones that satisfy our constraints, termed as valid step candidates.
+Among the valid candidates, DeepEdit takes the most important one as the new reasoning step for further reasoning, which is the (new) knowledge that is most helpful to answer the target question.
+Our depth-first search strategy efficiently increases the reasoning depth and enhances LLMs to produce the coherent reasoning chains involving new knowledge.
+The advantages of our DeepEdit on KE are three-folds.
+First, it does not need to retrain the LLMs like prior model editing methods Zhu et al.
+(2020), which is complex and expensive.
+Our DeepEdit works during inference with negligible costs.
+Second, our DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions.
+Third, our DeepEdit enhances LLMs to soundly utilize new knowledge to produce coherent reasoning chains in accordance with new knowledge.
+To provide more precise and challenging assessments of KE methods, we build two new benchmarks MQuAKE-2002, and MQuAKE-hard, which resolve the knowledge-conflicting annotation mistakes in the popular KE benchmark MQuAKE-3k Zhong et al.
 (2023).
-However, classical constrained decoding methods usually require access to the token-wise distribution, which is not provided by many black-box LLMs like ChatGPT OpenAI (2022).
-In our work, to apply our constraints to the black-box LLMs’ decoding, we breakthrough this restriction by changing the granularity of decoding from tokens to the reasoning steps.
-We design all decoding actions over the step level, which bypasses the requirement of accessing the token-wise distributions.
-At every iteration, we take one generated reasoning step and a few retrieved new facts as the step candidates.
-We verify the step candidates and filter out the ones that do not satisfy our constraints.
-Since there could be multiple candidates that satisfy all constraints every iteration.
-We operate the depth-first search that store all the valid step candidates in a stack and greedily search for the optimal path of LLMs’ multi-hop reasoning with new knowledge.
-We name our method as DeepEdit, the depth-first search-based decoding for knowledge editing.
-The advantages of our DeepEdit on KE are three-folds.
-First, it does not need to retrain the LLMs like the classical model editing methods Zhu et al.
-(2020), which is complex and expensive.
-Second, the new knowledge directly contributes to the LLMs’ decoding by acting as the reasoning step candidates.
-This facilitates more effective reasoning with new knowledge.
-Third, the influence of new knowledge is monitored with decoding constraints, leading to more controllable and traceable reasoning with new knowledge.
-Overall, our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness.
-DeepEdit can be flexibly applied to all the black-box LLMs.
 We evaluate DeepEdit on the KE benchmarks MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al.
@@
 (2022a).
-The latter two are new benchmarks that we build to provide more precise and challenging assessments of KE methods.
-Qualitatively, DeepEdit leads to more succinct and faithful reasoning of LLMs that soundly incorporate new knoweldge, as shown in Figure 8.
-We also observe substantial quantitative improvements in the KE question-answering accuracy achieved by DeepEdit.
-Overall, DeepEdit improves the KE for the strong LLMs text-davinci-003, gpt-3.5-turbo-instruct, and llama2 by a significant margin OpenAI (2022); Touvron et al.
+Qualitatively, DeepEdit enhances LLMs to produce more succinct and faithful reasoning with new knowledge, as shown in Figure 9.
+Quantitatively, DeepEdit improves the KE for popular LLMs by a significant margin OpenAI (2022); Touvron et al.
 (2023).
 2 Related Work
-Constrained Decoding
-Constrained decoding Qin et al.
-(2020, 2022); Lu et al.
-(2020, 2021); Kumar et al.
-(2022); Liu et al.
-(2023); Geng et al.
-(2023) is a classical natural language processing topic, which is mainly applied to control the lexicon, sentiment or style of the generated text.
-FUDGE Yang and Klein (2021) employs weighted decoding (WD) by directly adjusting the probabilities of the vocabulary distribution with an auxiliary classifier probability.
-NADO Meng et al.
-(2022c) suggests sampling from a vocabulary distribution similar to an approximation of the sequence-level oracle.
-GeDi Krause et al.
-(2020) and DExperts Liu et al.
-(2021) also adopt the weighted decoding approach through the training of generative classifiers.
-Constrained decoding methods generally require access to the LLMs’ token-wise distribution.
-Different from them, our DeepEdit avoids the requirements of distribution access by decoding at the reasoning step level.
-We extend the applications of constrained decoding to KE.
-We design the semantic constraints to soundly incorporate new knowledge in LLMs’ reasoning, and verify the complex semantic constraints with LLMs.
-Figure 3: Comparison of different KE methods.
-(upper) Model editing was the most popular KE paradigm, which edits the model parameters to update knowledge.
-(middle) Recently, in-context editing has become popular, which relies on the LLMs’ capability of utilizing new knowledge in the prompts.
-(lower) We propose DeepEdit for KE, which directly controls the LLMs’ decoding to follow new knowledge.
 Knowledge Editing
@@
 Different from them, we propose the new decoding method DeepEdit for KE, which effectively augments LLMs on KE during the decoding stage.
-We illustrate a comparison between DeepEdit and existing KE methods in Figure 3.
+We illustrate a comparison between DeepEdit and existing KE methods in Table 1.
 Specifically, in our work, we consider a challenging setting of knowledge editing: the multi-hop question answering (QA) with knowledge editing Zhong et al.
@@
 Our work improves multi-hop QA with KE by applying constraints to the reasoning of LLMs, which directly controls LLMs’ outputs to follow new knowledge.
-Figure 4: Comparison between classical constrained decoding (at the token level) (upper) and our decoding at the reasoning step level (lower).
-Differences: 1) The token candidates is the predefined vocabulary, while the step candidates include a generated step and the retrieved facts.
-2) The token candidates are ranked by generators’ outputs, while the step candidates are filtered by the constraints.
-Self-refinement of Large Language Models
-Some work has explored to use LLMs to revise the output of themselves in dialogue response, sentiment reversal, and mathematical reasoning Madaan et al.
-(2023); Shridhar et al.
-(2023b, a); Welleck et al.
-(2022); Zheng et al.
-(2023b); Madaan et al.
+Figure 3: When producing every reasoning step, we retrieve the edited facts that are close to the generated step in the semantic space (left).
+Only the edited facts that satisfy all the constraints, i.e., valid edited facts, are considered as the step candidates (right).
+KE Method Editing Editing Editing Model Parameters Input Prompts Outputs MEND Mitchell et al. (2021) MEMIT Meng et al. (2022c) ICE Cohen et al. (2023) MeLLo Zhong et al. (2023) DeepEdit (Ours)
+Table 1: Comparison between KE methods.
+Constrained Decoding
+Constrained decoding Qin et al.
+(2020, 2022); Lu et al.
+(2020, 2021); Kumar et al.
+(2022); Liu et al.
+(2023); Geng et al.
+(2023) is a classical natural language processing topic, which is mainly applied to control the lexicon, sentiment or style of the generated text.
+FUDGE Yang and Klein (2021) employs weighted decoding (WD) by directly adjusting the probabilities of the vocabulary distribution with an auxiliary classifier probability.
+NADO Meng et al.
+(2022c) suggests sampling from a vocabulary distribution similar to an approximation of the sequence-level oracle.
+GeDi Krause et al.
+(2020) and DExperts Liu et al.
+(2021) also adopt the weighted decoding approach through the training of generative classifiers.
+Constrained decoding methods generally require access to the LLMs’ token-wise distribution.
+Different from them, our DeepEdit avoids the requirements of distribution access by decoding at the reasoning step level.
+We extend the applications of constrained decoding to KE.
+We design the semantic constraints to soundly incorporate new knowledge in LLMs’ reasoning, and verify the complex semantic constraints with LLMs.
+3 Methodology
+Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) of large language models (LLMs) Zhong et al.
 (2023).
-However, it is difficult to rely on LLMs to accommodate the new knowledge by themselves in the self-refinement, since the new knowledge is generally conflicted with the parametric knowledge memorized by LLMs.
-Different from them, we do not rely on LLMs to revise the outputs by themselves.
-We directly control the decoding to soundly incorporate new knowledge with decoding constraints.
-We modify the LLMs’ outputs using knowledge retrieval and depth-first search to satisfy the constraints.
-We hope that our work demonstrates more potential to conduct KE for LLMs in the decoding stage.
-3 Methodology
-Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) Zhong et al.
-(2023), where LLMs should answer the questions corresponding to a chain of facts including new knowledge.
-This task assumes a database storing all new facts that LLMs should follow, and only a few facts can influence the ground-truth answer to a specific question.
-Our KE idea is to first design the decoding constraints that can regulate LLMs to soundly incorporate new knowledge in reasoning and then enforce LLMs’ decoding to follow the constraints.
-In this section, we will introduce how we incorporate our idea into the design of a new KE method DeepEdit (Depth-first Search based Decoding for Knowledge Editing) to improve the KE of black-box LLMs.
-3.1 Decoding at the Reasoning Step Level
-Many classical constrained decoding methods Zhang et al.
-(2023a); Meng et al.
-(2022c) use an auxiliary model to adjust the token-wise distribution during decoding.
-They manipulate the distribution before every token’s generation to satisfy the constraints.
-Therefore, they require accessibility to the LLMs’ output distribution, which is not feasible for many black-box LLMs.
-In this work, we change the granularity of decoding from the token level to the reasoning step level.
-Our decoding is designed at the step level: instead of taking the tokens predefined in a fixed vocabulary as the candidates, we take the reasoning steps as the candidates.
-At every iteration, we let LLMs to produce one reasoning step and retrieve a fixed number of new facts in parallel to the reasoning step as the step candidates.
-The new fact retrieval is based on the nearest neighbor search with the sentence-level embeddings given by BERT.
-The motivation of this design is that the new facts that are semantically close to the generated facts are more likely to satisfy the constraints.
-Taking the single generated step and the new facts as the step candidates, we filter a valid candidate and take it as the new reasoning step.
-Note that there may exist more than one candidates that satisfy the constraints.
-We will introduce how we design the decoding constraints for KE in Section 3.2, how we verify the constraints in Section 3.3 and Section 3.4, and how we process the multiple valid step candidates to search for the optimal path of LLMs’ multi-hop reasoning in Section 3.5.
-Our decoding design bypasses the requirement of accessing the token-wise distributions, since all the decoding actions are defined over the step level.
-Our decoding strategy is free to be applied to black-box LLMs in order to produce satisfactory outputs in accordance with new knowledge.
-The details are introduced next.
-Figure 5: (left) At every decoding iteration, we store the reasoning step candidates (including the generated step and the retrieved knowledge) that satisfy the local constraints in a stack.
-(right) The global verifier determines the feasibility of the whole reasoning process by checking whether the answer is found.
-If not feasible, we pop the status stack to use another reasoning subprocess that satisfies the local verifier.
-3.2 Designing Decoding Constraints for Knowledge Editing
-To correctly answer multi-hop questions in accordance with new knowledge Zhong et al.
-(2023), LLMs need to generate a sequence of reasoning steps to find the answer.
-We propose the following constraints for the LLMs’ reasoning with KE.
+This task considers the questions that correspond to a chain of facts including new knowledge.
+It assumes a database that stores all new facts, which LLMs should adhere to, and posits that only a few facts can influence the ground-truth answer to each question.
+In this section, we will introduce how we integrate the decoding constraints into the design of a novel KE method, DeepEdit (Depth-first Search based Decoding for Knowledge Editing), to enhance the KE of black-box LLMs.
+3.1 Constraint Designing for Knowledge Editing
+On a LLM’s multi-step reasoning chain to answer a multi-hop question involving new knowledge, we find the following constraints on reasoning steps can help LLMs to give the correct answer:
   1. 1.
-    [Conciseness] Every reasoning step should be different from the prior ones.
+    Conciseness: Every step is identical.
   2. 2.
-    [Coherence] Every reasoning step should be relevant to the last reasoning step.
+    Coherence: Adjacent steps are coherent.
   3. 3.
-    [Receptiveness] Every reasoning step cannot contradict to any new knowledge.
+    Receptiveness: No conflicts exist between reasoning steps and new knowledge.
   4. 4.
-    [Pertinence] Every reasoning step should be relevant and helpful to finding the answer.
-The above constraints are not automatically satisfied by LLMs when reasoning with new knowledge.
-However, they are essential to helping LLMs to find the correct answers in accordance with new knowledge.
-Therefore, in our work, we propose a new decoding method to enforce the LLMs to follow the above constraints so as to improve the KE performance.
-We first analyze the constraints in detail.
-Among the above four constraints, [Conciseness] guarantees that the LLMs’ reasoning is not in a loop.
-[Coherence] guarantees the coherence of reasoning, while [Pertinence] guarantees the relevance of every reasoning step to the question.
-[Receptiveness] ensures that LLMs can utilize new knowledge for reasoning.
+    Pertinence: Every reasoning step should be relevant to the target question.
+Among the above four constraints, Conciseness removes loops in the reasoning chain; Coherence guarantees the coherence of reasoning; Receptiveness ensures the LLMs’ receptiveness to new knowledge; Pertinence guides LLMs to answer the target question.
 We visualize different constraints in Figure 2.
-Among the constraints, we regard [Conciseness], [Coherence], and [Receptiveness] as local ones: we can verify whether they are satisfied given only partial reasoning steps.
-The [Pertinence] constraint is regarded as a global constraint: we can verify it based on a complete reasoning process.
-By following the above constraints, LLMs can orderly place new knowledge at appropriate positions in the reasoning chain and avoid the memorized knowledge that is conflicted with new knoweledge.
-Next, we will introduce how we verify the above constraints and control the LLMs’ decoding to satisfy the constraints.
-3.3 Verifier for Local Constraints
-To verify the [Conciseness] constraint, by default, whenever a new reasoning step is given, we may compare it with all the prior reasoning steps.
-However, this is no doubt an expensive solution at the cost of $\mathcal{O}(r^{2})$ times of verification given as many as $r$ reasoning steps.
-We observe that the repeated reasoning step generally occurs in the adjacent steps.
-Therefore, to reduce the computational costs, we propose an approximate verification method to verify the constraint of [Conciseness].
-Specifically, for every reasoning step, we only compare it with the last reasoning step, so that the whole time of doing the verification is $\mathcal{O}(r)$.
-Figure 6: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k.
-The chosen step candidate in every iteration that satisfies all the constraints are in green.
-The dropped step candidates are followed by the constraints that they do not satisfy.
-To satisfy the [Coherence] constraint, we need to evaluate whether the current reasoning step is relevant to the last reasoning step.
-However, it is challenging to directly and precisely evaluate the relevance through LLMs.
-We use the LLM based verifiers equipped with a hand-crafted prompt to verify whether the adjacent reasoning steps are semantically equivalent or related to verify the constraints of [Conciseness] and [Coherence] respectively.
-Verifying the [Receptiveness] constraint, by default, requires determining whether the contradiction exists in the pair of the reasoning step and all the items in the basis of new knowledge.
-The time complexity is $\mathcal{O}(rk)$, where $k$ is the number of new knowledge items and a large value in practice.
-Therefore, to reduce the time complexity, we retrieve the relevant knowledge items that are semantically closest to the reasoning step and then assess the contradictions.
-The idea is that if two sentences are contradictory to each other, they must first be semantically relevant.
-In this way, we reduce the time complexity from $\mathcal{O}(rk)$ to $\mathcal{O}(r\log(k))$.
-Also, we design a LLM-based verifier with a hand-crafted prompt to assess whether the retrieved facts are contradictory to the generated reasoning step.
-3.4 Verifier for Global Constraints
-Whether a reasoning step is helpful in finding the final answer is hard to justify until the final answer is found.
-Therefore, we only verify the [Pertinence] constraint globally for each reasoning step given the full completed reasoning process.
-When the intermediate reasoning steps are not helpful but even barrier the final answer, the reasoning likely takes more steps to find the final answer.
-Based on this prior knowledge, we set the maximum reasoning steps as $r$, so that the reasoning is terminated when an answer is not found given more than $r$ steps.
-We set $r=5$ by default in our experiments.
-3.5 Decoding with Depth-first Search
-As we have introduced in Section 3.1, there may exist multiple step candidates that satisfy all the constraints. In this case, how can we decide which step is the next step?
-To answer this question, let’s first recall how LLMs how to decide which is next token.
-The most simple method is the greedy decoding, which just takes the highest logit token candiate as the next token.
-In our case, since we do not have scores to rank step candidates but only the discrete constraint verifications to filter the candidates, we cannot simply take the “best” step as the next reasoning step.
-Can we have other alternate solutions?
-Let’s see a more advanced token-level decoding method: beam search.
-Beam search takes multiple candidate tokens as the next tokens and store them in a stack to do a best-first search in the token sequence generation.
-We borrow the idea of the beam search to process the multiple valid step candidates with the depth-first search.
-Specifically, we store all the step candidates that satisfy the constraints in a stack, as shown in Figure 5.
-Then, in every iteration, we pop the reasoning step stored at the head of the stack as the next reasoning step and do further reasoning based on it.
-We name our method as DeepEdit: depth-first search based decoding for knowledge editing.
-We visualize our DeepEdit in Figure 1(c).
-Also, we present the iteration-wise outputs of our DeepEdit on an instance in Figure 6.
+Intuitively, the above constraints are easy to follow for LLMs when there is no new knowledge.
+However, the LLMs’ utilization of new knowledge is affected by the hallucination, which makes the above constraints easy to break when incorporating new knowledge.
+In this work, we apply the above constraints to incorporate the new knowledge into LLMs’ decoding.
+3.2 Constraint Verification on Every Step
+We term the reasoning steps that satisfy all the constraints as valid steps.
+Because an invalid reasoning step could mislead LLMs into generating incorrect subsequent steps, we verify the constraints during the decoding of each reasoning step.
+At every iteration, we take both the parametric and new knowledge as the step candidates.
+First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be useful to answer the target question.
+Second, we retrieve $N$ new knowledge closest to the generated reasoning step.
+The distance between knowledge is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al.
+(2019).
+The closer the context of the new knowledge is to the generated step, the more likely it is that the new knowledge will be helpful in answering the target question.
+We take these $N+1$ knowledge as the reasoning step candidates, with examples shown in Figure 3.
+We verify the constraints on the above step candidates to find the valid step candidates.
+In principle, every constraint’s verification can be seen as a binary classification problem on a pair of sentences.
+For example, Coherence’s verification can be completed by assessing whether a pair of reasoning steps is relevant to each other or not.
+Therefore, one option to do constraint verification is to train a binary classifier for every constraint.
+However, such a solution is not data or time-efficient.
+In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers.
+Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from the instances of the MQuAKE dataset Zhong et al.
+(2023).
+In our experiments, we set $D=2$ and $N=2$ by default.
+3.3 Efficient Reasoning with Depth-first Search
+At every iteration, there may exist more than valid step candidates.
+We can take every one of them as the new reasoning step in parallel and make further reasoning based on each of them separately.
+This design can be seen as a breadth-first search of the LLMs’ multi-hop reasoning chain incorporating new knowledge.
+It is undoubtedly time-consuming because the time complexity grows exponentially with the number of reasoning steps.
+This raises a natural question:
+Can we more efficiently expand the depth of reasoning chains during decoding?
+Our answer is yes.
+We propose to take the most important step candidate as the reasoning step for further reasoning.
+Actually, different valid step candidates are not equally important.
+First, the edited facts should be more important than the generated step, since the former generally contains new knowledge unknown by LLMs.
+Second, the edited facts that are closer to the generated step are more important, which are more helpful in answering the target question due to their relevance to the LLM’s generated reasoning context.
+Based on the above prior, in every iteration, we select the most important knowledge as the next reasoning step and do further reasoning based on it, as shown in Figure 4.
+Figure 4: At every iteration, we verify the decoding constraints over parametric and new knowledge to find the valid step candidates.
+Then we take the most important candidate as the next reasoning step to efficiently increase the reasoning depth.
+We name our method as DeepEdit: depth-first search-based constrained decoding for knowledge editing.
+DeepEdit takes the most important step candidate as the new reasoning step at every iteration and efficiently produces the multi-hop reasoning incorporating new knowledge.
+We summarize DeepEdit in Algorithm 1 and visualize it in Figure 1.
+Furthermore, we present the iteration-wise outputs of our DeepEdit in Figure 8.
+Assume the number of reasoning hops as $h$, and the number of valid step candidates per iteration as $c$.
+The time complexity of breadth-first search is $c^{h}$.
+In contrast, our DeepEdit has a much lower time complexity of $\mathcal{O}(h)$ thanks to our depth-first search design considering the importance of different candidates.
+Empirical results validate DeepEdit’s superior efficiency compared with the breadth-first search.
 The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al.
 (2023).
-Different from them, our DeepEdit’s depth-first search is to enforce the LLMs’ decoding to satisfy the constraints derived for KE from a global perspective.
-4 New Benchmarks for More Precise and Challenging Evaluation
+Different from them, our DeepEdit’s depth-first search is designed to efficiently find the multi-hop reasoning chain incorporating new knowledge.
+Our DeepEdit is free to be applied to all black-box LLMs: it bypasses the requirement of accessing the token-wise distributions because all the decoding actions are defined at the step level.
+Algorithm 1 DeepEdit: Depth-first Search based Constrained Decoding for Knowledge Editing.
+Input: The query question $q$. The prompt $p$. The set of new knowledge $\mathcal{K}$. The number of retrieved new knowledge $N$. Maximum number of reasoning depth $r$.
+Output: The answer $a$ to the query question.
+1:for $i\leftarrow 1$ to $r$ do
+2:     $o\leftarrow\mbox{LLM}(p)$
+3:     if $o$ is the answer then
+4:         return $o$
+5:     $\hat{\mathcal{K}}\leftarrow\mbox{search}(o,\mathcal{K},n)$triangleright retrieve the $n$ semantically similar new knowledge to $o$, ranked by the distance to $o$ in the ascending order.
+6:     for $j\leftarrow 1$ to $n$ do
+7:         $k\leftarrow\hat{\mathcal{K}}[i]$
+8:         $nk\leftarrow\mbox{False}$triangleright $nk$ is the boolean variable to decide whether some new knowledge is valid.
+9:         if ($k$ satisfies all the constraints) then
+10:              $nk\leftarrow\mbox{True}$
+11:              $p\leftarrow p+k$
+12:              break
+13:     if not $nk$ then
+14:         $p\leftarrow p+o$      return $o$
+4 New Benchmarks for More Precise and Challenging Evaluation of KE
 We provide two new benchmarks for the evaluation of KE methods: MQuAKE-2002 and MQuAKE-hard.
-Both are built from the MQuAKE Zhong et al.
-(2023), a recent KE benchmark that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge.
-Every instance in MQuAKE includes a multi-hop question and one to four related edited facts, every of which can alter the ground-truth answer to the question.
+Both are built based on MQuAKE Zhong et al.
+(2023), a recent KE dataset that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge.
+Every instance in MQuAKE includes a multi-hop question and several edited facts, every of which can alter the ground-truth answer to the question.
+4.1 Annotation Mistakes in One-Third Instances of MQuAKE
 Zhong et al.
-(2023) suggests using a randomly sampled 3,000 instance subset of MQuAKE to do the evaluation, which reduces the experimental costs.
-This evaluation benchmark is termed as MQuAKE-3k.
-4.1 Annotation Mistakes in One-Third Instances of MQuAKE-3k
+(2023) suggests using a randomly sampled 3,000 instance subset of MQuAKE to do the evaluation, which reduces the experimental costs, known as MQuAKE-3k.
 There are two issues of using MQuAKE-3k to evaluate KE.
-The first issue is that the new knowledge from different instances MQuAKE-3k can cause conflicts and mistakes to the ground-truth answers.
+The first issue is that the new knowledge from different instances MQuAKE-3k can cause conflicts and mistakes in the ground-truth answers.
 In other words, the ground-truth answer from instance $A$ can be altered by the new knowledge from another instance $B$.
 We show an example of knowledge conflicts in MQuAKE-3k in Figure 5.
-These knowledge conflicts will make the ground-truth answers incorrect given the new knowledge that are conflicted with the answers because the inference on every instance would retrieve the new knowledge from all instances.
-We observe that 998 instances’ ground-truth answers are broken by the new knowledge from other instances.
+These knowledge conflicts will make the ground-truth answers incorrect given the conflicted new knowledge because the inference on every instance could retrieve the new knowledge from all instances.
+We observe that 998 instances’ ground-truth answers are broken by the new knowledge conflicts across instances.
 4.2 New Benchmark MQuAKE-2002 for More Precise Evaluation
-To address the issue of annotation mistakes in MQuAKE-3k, we provide a new subset of MQuAKE-3k, which does not have any knowledge conflict across instances.
-This subset includes 2,002 instances, so we term it as MQuAKE-2002.
+To address the issue of annotation mistakes in MQuAKE-3k, we provide a new KE benchmark based on MQuAKE, which does not have any knowledge conflict across instances.
+This benchmark includes 2,002 instances, so we term it as MQuAKE-2002.
 We filter out the instances of which the ground-truth answers are broken by the new knowledge from other instances to produce MQuAKE-2002.
-Compared with MQuAKE-3k, our MQuAKE-2002 provides a more precise evaluation for knowledge editing methods, since it remove the annotation mistakes due to knowledge conflicts across instances.
+Compared with MQuAKE-3k, our MQuAKE-2002 provides a more precise evaluation for KE methods, since it removes the annotation mistakes due to knowledge conflicts across instances.
 The data statistics of MQuAKE-2002 is provided in Table 2.
-4.3 New Benchmark MQuAKE-hard for More Challenging Evaluation
-The second issue of MQuAKE-3k is that more than 66% instances in MQuAKE-3k only contain at most two edited facts that influence the answers, which are not challenging enough to evaluate the knowledge editing methods on handling multiple edited facts that can alter the ground-truth reasons.
-we construct a more challenging subset of MQuAKE by selecting the instances that contain the highest number of edited facts per instance.
-We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts.
-MQuAKE-hard has no overlap with MQuAKE-3k.
-We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts.
-MQuAKE-hard has no overlap with MQuAKE-3k.
+4.3 New Benchmark MQuAKE-hard for Challenging Evaluation
+The second issue of MQuAKE-3k is that more than 60% instances in it only contain at most two edited facts that influence the answers, which are not challenging enough to evaluate the KE methods on handling multiple edited facts that can alter the ground-truth answers.
+We construct a more challenging subset of MQuAKE by selecting the instances that contain the highest number of edited facts per instance excluding MQuAKE-3k.
+We term this challenging set as MQuAKE-hard, which includes 429 instances and every instance contains four edited facts that can change the ground-truth answer.
 The data statistics of MQuAKE-hard is also provided in Table 2.
-Benchmark # Instances # Hops per Instance # Edited Facts per Instance # Conflicted Instances MQuAKE-3k Zhong et al. (2023) 3,000 3.0 2.0 998 MQuAKE-2002 2,002 2.7 2.2 0 MQuAKE-hard 4,29 4.0 4.0 0
+Benchmark # Instances # Hops per Instance # Edited Facts per Instance # Conflicted Instances MQuAKE-3k Zhong et al.
+(2023) 3,000 3.0 2.0 998 MQuAKE-2002 (Ours) 2,002 2.7 2.2 0 MQuAKE-hard (Ours) 429 4.0 4.0 0
 Table 2: Data statistics of different benchmarks.
@@
 5 Experiments
-In this section, we evaluate the KE performance of our DeepEdit method when applied to LLMs, and compare against a strong baseline MeLLo Zhong et al.
+In this section, we evaluate the KE performance of our DeepEdit method when applied to LLMs, and compare against strong KE baselines Zhong et al.
 (2023).
@@
 (2023), MQuAKE-2002, and MQuAKE-hard for evaluation.
-The introduction of the above benchmarks are provided in Section 4.
 We follow the existing work Zhong et al.
@@
 (2023) to set the KE batch size as 100 and full batch for evaluation.
-The KE batch size means the number of instances that provide new knowledge for retrieval.
+The KE batch size means the number of instances that provide new knowledge for retrieval when answering every question.
 Compared methods.
@@
 (2022b), IKE Zheng et al.
-(2023a), ICE Cohen et al.
+(2023), ICE Cohen et al.
 (2023), and MeLLo Zhong et al.
@@
 The former two KE methods are only applicable to white-box LLMs, while the latter three can fit the block-box LLMs.
-Method MQuAKE-3k MQuAKE-hard batch size of 100 full batch batch size of 100 full batch llama2-7b w/ MEND Mitchell et al.
-(2021) 7.2 3.9 0.5 0.2 llama2-7b w/ MEMIT Meng et al.
-(2022b) 7.9 4.2 0.9 0.5 llama2-7b w/ IKE Zheng et al.
-(2023a) 12.5 6.2 1.2 0.5 llama2-7b w/ ICE Cohen et al.
-(2023) 12.9 6.3 1.4 0.7 llama2-7b w/ MeLLo Zhong et al.
-(2023) 16.3 10.8 1.9 1.6 llama2-7b w/ DeepEdit (Ours) 27.6 11.2 14.5 7.0 gpt-3.5 w/ IKE Zheng et al.
-(2023a) 15.5 8.3 1.6 1.2 gpt-3.5 w/ ICE Cohen et al.
-(2023) 16.1 8.5 2.1 1.4 gpt-3.5 w/ MeLLo Zhong et al.
-(2023) 29.9 20.0 3.7 1.6 gpt-3.5 w/ DeepEdit (Ours) 47.2 38.0 51.3 48.0 text-davinci-003 w/ IKE Zheng et al.
-(2023a) 18.2 8.9 1.9 1.2 text-davinci-003 w/ ICE Cohen et al.
-(2023) 19.0 8.6 2.3 1.2 text-davinci-003 w/ MeLLo Zhong et al.
-(2023) 32.2 25.6 4.7 2.3 text-davinci-003 w/ DeepEdit (Ours) 49.1 38.8 52.0 49.2
+Method MQuAKE-3k MQuAKE-hard MQuAKE-2002 100 edited all edited 100 edited all edited 100 edited all edited LLaMA2-7B-Chat w/ MEND Mitchell et al.
+(2021) 7.2 3.9 0.5 0.2 7.7 4.1 LLaMA2-7B-Chat w/ MEMIT Meng et al.
+(2022b) 7.9 4.2 0.9 0.5 8.3 4.8 LLaMA2-7B-Chat w/ IKE Zheng et al.
+(2023) 12.5 6.2 1.2 0.5 12.9 6.5 LLaMA2-7B-Chat w/ ICE Cohen et al.
+(2023) 12.9 6.3 1.4 0.7 13.2 6.5 LLaMA2-7B-Chat w/ MeLLo Zhong et al.
+(2023) 16.3 10.8 1.9 1.6 17.1 11.8 LLaMA2-7B-Chat w/ DeepEdit (Ours) 27.6 11.2 14.5 7.0 29.5 12.9 GPT-3.5-Turbo-Instruct w/ IKE Zheng et al.
+(2023) 15.5 8.3 1.6 1.2 17.8 11.0 GPT-3.5-Turbo-Instruct w/ ICE Cohen et al.
+(2023) 16.1 8.5 2.1 1.4 18.1 11.5 GPT-3.5-Turbo-Instruct w/ MeLLo Zhong et al.
+(2023) 29.9 20.0 3.7 1.6 31.2 25.1 GPT-3.5-Turbo-Instruct w/ DeepEdit (Ours) 47.2 38.0 51.3 48.0 60.1 53.7 Text-Davinci-003 w/ IKE Zheng et al.
+(2023) 18.2 8.9 1.9 1.2 19.1 10.2 Text-Davinci-003 w/ ICE Cohen et al.
+(2023) 19.0 8.6 2.3 1.2 19.5 10.8 Text-Davinci-003 w/ MeLLo Zhong et al.
+(2023) 32.2 25.6 4.7 2.3 33.5 27.8 Text-Davinci-003 w/ DeepEdit (Ours) 49.1 38.8 52.0 49.2 60.9 54.8
 Table 3: Experimental results (accuracy; %) on the dataset MQuAKE-3k and MQuAKE-hard.
-The batch size means the number of instances providing the edited facts for knowledge retrieval.
-gpt-3.5 stands for the model gpt-3.5-turbo-instruct.
+“$k$ edited” denotes the KE batch size, which means the number of instances providing new knowledge on answering every question.
+Models with are closed source.
 The best KE result on every LLM is highlighted in bold font.
-Method MQuAKE-2002 gpt-3.5 w/ MeLLo 25.1 gpt-3.5 w/ DeepEdit 53.7
-Table 3: Experimental results (accuracy; %) of the full-batch knowledge editing on the benchmark MQuAKE-2002.
-Verifier [Conciseness] [Coherence] [Receptiveness] llama2-7b 82.3 77.9 86.3 gpt-3.5 92.8 91.5 92.2
-Table 4: Experimental results (verification accuracy; %) of our verifiers for different constraints on the dataset MQuAKE-3k.
-gpt-3.5 w/ DeepEdit 49.1 $\Delta$ w/o [Conciseness] 46.1 $\downarrow 6\%$ w/o [Coherence] 42.9 $\downarrow 13\%$ w/o [Receptiveness] 45.2 $\downarrow 8\%$ w/o [Pertinence] 46.6 $\downarrow 5\%$
-Table 5: Experimental results (accuracy; %) on the dataset MQuAKE-3k of DeepEdit with KE batch size of 100.
+LLM as Constraint Verfier Conciseness Coherence Receptiveness Pertinence LLaMA2-7B-Chat 82.3 77.9 86.3 78.2 GPT-3.5-Turbo-Instruct 92.8 91.5 92.2 86.7
+Table 4: Experimental results (verification accuracy; %) of our verifiers for different constraints.
 Model configuration.
 We set the hyper-parameters of the baseline methods as suggested by their papers.
-We apply the KE methods to the strong LLM text-davinci-003, gpt-3.5-turbo-instruct (short as gpt-3.5), and llama2-7b (short as llama2) OpenAI (2022); Touvron et al.
+We apply the KE methods to the popular LLMs GPT-3.5-Turbo-Instruct, Text-Davinci-003, and LLaMA2-7B-Chat OpenAI (2022); Touvron et al.
 (2023) for evaluation.
@@
 By default, we set the decoding temperature as 0.0 to minimize the randomness of LLMs’ outputs.
-The DeepEdit’s verifiers use the same LLM as the generator.
-Figure 8: A case study for MeLLo and our DeepEdit on the dataset MQuAKE-3k. The new knowledge are in green. We report the prediction of text-davinci-003 with MeLLo and our DeepEdit.
+Our constraint verifiers use the same LLM as the generator.
 5.2 Overall Performance
 We incorporate DeepEdit with several choices of LLMs to test its effectiveness on editing LLMs’ knowledge.
-We report the test accuracy on MQuAKE-3k and MQuAKE-hard in Table 2.
-When the KE batch size is 100, our DeepEdit improves the accuracy of question answering with KE by 69%, 58%, and 52% over MeLLo on the LLMs llama2, gpt-3.5-turbo-instruct, and text-davinci-003 respectively over the strong baseline method MeLLo on MQuAKE-3k.
-On MQuAKE-hard, our DeepEdit outperforms the strong baseline method MeLLo by $\times$7, $\times$12, and $\times$10 times in terms of question answering accuracy with KE on the LLMs llama2, gpt-3.5, and text-davinci-003 respectively.
-The experimental results on MQuAKE-2002 is shown in Table 3.
-When the KE batch size is full batch, our DeepEdit still outperforms the baseline methods and improves the KE accuracy by a significant margin.
+We report the test accuracy on MQuAKE, MQuAKE-3k, MQuAKE-2002 in Table 3.
+When the KE batch size is 100, our DeepEdit improves the accuracy of question answering with KE by 69%, 58%, and 52% over MeLLo on the GPT-3.5-Turbo-Instruct, Text-Davinci-003, LLaMA2-7B-Chat respectively over the strong baseline method MeLLo on MQuAKE.
+On MQuAKE-hard, our DeepEdit outperforms the strong baseline method MeLLo by $\times$7, $\times$12, and $\times$10 times in terms of question answering accuracy with new knowledge on the LLMs LLaMA2-7B-Chat, Text-Davinci-003, and GPT-3.5-Turbo-Instruct respectively.
+When the KE batch is full batch, our DeepEdit still outperforms the baseline methods and improves the KE accuracy by a significant margin.
 Our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness.
-The above advantages are especially important when the LLMs have to utilize multiple new knowledge to get the correct answers, since more new knowledge requires the LLMs to be correctly aware of the new knowledge on more reasoning steps.
+The above advantages are especially significant when the LLMs have to utilize multiple new knowledge to get the correct answers because more new knowledge requires LLMs to be correctly aware of the new knowledge on more reasoning steps.
 The correct knowledge incorporation with our decoding constraints significantly improves the KE performance given more new knowledge.
-Overall, our DeepEdit achieves substantial improvements on KE for LLMs and consistently outperforms the baseline methods on different benchmarks.
-5.3 Analysis on Verifiers
-We conduct deep analysis on the different verifiers of DeepEdit.
-We first test the accuracy of different verifiers to verify the corresponding constraint.
-We randomly sample 4,389 pairs of reasoning steps from the ground-truth reasoning steps of MQuAKE as the positive pairs that satisfy the constraints and randomly pairing the same number of steps from different questions as the negative examples.
-We evaluate different verifiers based on llama2 and gpt-3.5 on the sampled reasoning step pairs.
+Overall, our DeepEdit achieves substantial improvements on KE for LLMs and consistently outperforms the baseline KE methods on various benchmarks.
+GPT-3.5-Turbo-Instruct w/ DeepEdit 49.1 $\Delta$ w/o Conciseness 46.1 $\downarrow 6\%$ w/o Coherence 42.9 $\downarrow 13\%$ w/o Receptiveness 45.2 $\downarrow 8\%$ w/o Pertinence 46.6 $\downarrow 5\%$
+Table 5: Experimental results (accuracy; %) on the dataset MQuAKE of DeepEdit with KE batch size of 100.
+5.3 Accuracy of Constraint Verification
+We conduct an in-depth analysis on the constraint verification of our constraints.
+We randomly sample 4,000 pairs of sentences from MQuAKE that satisfy our constraints as positive samples and the same number of sentence pairs from different instances as the negative samples.
+We evaluate different verifiers based on LLaMA2-7B-Chat and GPT-3.5-Turbo-Instruct on the above samples.
 We report the verification accuracy in Table 4.
-We observe high verification accuracy of around 80% and over 90% on different constraints achieved by llama-2 and gpt-3.5 respectively.
-Also, we observe that the performance of verifiers based on llama2 and gpt-3.5 are positively related to the LLMs’ performance with our DeepEdit, since the effective constraint assessment forms the basis of effective reasoning of LLMs with our DeepEdit.
+We observe high verification accuracy of around 80% and over 90% on different constraints achieved by LLaMA2-7B-Chat and GPT-3.5-Turbo-Instruct respectively.
+This result matches the performance gap between two LLMs on general reasoning tasks OpenAI (2022); Touvron et al.
+(2023).
+5.4 Efficiency of Depth-first Search
+As analyzed in Section 3.3, we utilize the depth-first search to efficiently increase the LLMs’ reasoning depth with new knowledge.
+We evaluate the efficiency of our depth-first search using the 3-hop and 4-hop questions in the MQuAKE-3k benchmark, which holds the longest reasoning chains among the instances.
+The methods we evaluate include: MeLLo, the breadth-first search-based decoding method introduced in Sec.
+3.3, and our depth-first earch.
+The average number of reasoning steps and the time of text generation on every instance taken by different methods are reported in Figure 6.
+where the generation time is on a Linux Server using three A6000 GPUs.
+We notice that, compared with the breadth-first search method, our DeepEdit significantly reduces the number of reasoning steps and the generation time.
+Our DeepEdit efficiently finds the reasoning chain with new knowledge by considering the importance of different step candidates.
+These results agree with the theoretical analysis in Sec.
+3.3, since our DeepEdit’ time complexity is $\mathcal{O}(h)$ in the base case, while the time complexity of the breadth-first search is $\mathcal{O}(c^{h})$, much higher than our DeepEdit.
+Compared with the baseline MeLLo, our DeepEdit leads to much less generation time because our output text is purely the reasoning steps without the redundant subquestions and edited facts kept by MeLLo.
+Figure 6: Inference costs of the baseline method MeLLo Zhong et al. (2023), breadth-first search introduced in Sec. 3.3, and our depth-first search.
 5.5 Ablation Study
@@
 The above results also validate that the LLMs’ outputs cannot automatically satisfy the constraints that represent the desired properties without our DeepEdit.
-5.5 Case Study
+More experiments and case studies can be found in the appendix.
+6 Conclusion
+We explore a new paradigm for knowledge editing of black-box LLMs: decoding with constraints.
+We have designed a new decoding method DeepEdit to improve the knowledge editing of LLMs.
+DeepEdit enhances LLM’s outputs to soundly and efficiently incorporate new knowledge in the multi-step reasoning.
+In addition to DeepEdit, we provide two new benchmarks: MQuAKE-2002 and MQuAKE-hard to provide more precise and challenging evaluations for knowledge editing methods.
+Extensive experiments demonstrate the effectiveness of our DeepEdit.
+Future work includes exploring other constraints that can improve knowledge editing for large language models.
+Acknowledgements
+This work was partially supported by a DARPA ANSR program FA8750-23-2-0004.
+The views and conclusions are those of the authors and should not reflect the official policy or position of DARPA or the U.S.
+Government.
+Appendix A Analysis on Knowledge Editing Batch Size
+We follow the existing work Zhong et al.
+(2023) to evaluate the influence of the KE batch sizes on the performance.
+A larger batch size leads to the more new knowledge collected from more instances for retrieval.
+We present the test accuracy on MQuAKE-3k with different batch sizes of KE in Figure 7.
+We observe that the performance of GPT-3.5-Turbo-Instruct with MeLLo and our DeepEdit decreases with a larger batch size of knowledge editing.
+The reason is that with a larger batch, the difficulty of retrieving the relevant new knowledge to the target questions becomes higher.
+On the other hand, our DeepEdit still exhibits significant and consistent improvements over the strong KE method MeLLo.
+Figure 7: Test accuracy (%) of question answering with new knowledge with different KE batch sizes. We present the performance of MeLLo and our DeepEdit applied on GPT-3.5-Turbo-Instruct.
+Figure 8: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k.
+The chosen step candidate in every iteration that satisfies all the constraints are in green.
+The dropped step candidates are followed by the constraints that they do not satisfy.
+Figure 9: A case study for MeLLo and our DeepEdit on the dataset MQuAKE-3k. The new knowledge are in green. We report the prediction of Text-Davinci-003 with MeLLo and our DeepEdit.
+Appendix B Analysis on the Data where the Edited Facts do not Change the Ground-Truth Answers
+Knowledge editing methods enhance the LLMs’ reasoning capability on the new knowledge.
+However, there exists the risk that the KE methods can degrade the LLMs’ performance on the data where no new knowledge changes their ground-truth answers.
+To evaluate whether our method DeepEdit degrades the performance of LLMs in this case.
+We do the following experiments, for every question-answering instance in MQuAKE-3k, we retrieve the edited facts only from other instances as the new knowledge, which would not change the ground-truth answer for the target question.
+In this setting, we first test the GPT-3.5-Turbo-Instruct without any KE methods, and compare the performance of GPT-3.5-Turbo-Instruct with the KE methods MeLLo and our DeepEdit.
+We present the performance in Table 6.
+Method MQuAKE-3k GPT-3.5-Turbo-Instruct 55.7 GPT-3.5-Turbo-Instruct w/ MeLLo Zhong et al. (2023) 54.2 GPT-3.5-Turbo-Instruct w/ DeepEdit (Ours) 55.5
+Table 6: Experimental results (accuracy; %) on the dataset MQuAKE-3k with no edited facts change the ground-truth answers.
+The empirical results show that our DeepEdit exhibits higher robustness to give the correct answers for the multi-hop question answering than MeLLo.
+The performance drop of GPT-3.5-Turbo-Instruct with our DeepEdit when no edited facts change the ground-truth labels is negligible.
+In addition to the above experiments, we also follow existing work Meng et al.
+(2022a); Zheng et al.
+(2023) to evaluate the Specificity scores of different knowledge editing methods on CounterFact given the KE batch size as full batch, which measures the accuracy of querying neighborhood prompts by Neighborhood Scores.
+In CounterFact, the neighborhood prompts share the same original object with the target prompt and these facts are not supposed to be edited.
+The results are presented in Table 7.
+Method CounterFact LLaMA2-7B-Chat 85.3 LLaMA2-7B-Chat w/ MEND Mitchell et al.
+(2021) 57.8 LLaMA2-7B-Chat w/ MEMIT Meng et al.
+(2022b) 80.9 LLaMA2-7B-Chat w/ MeLLo Zhong et al.
+(2023) 82.7 LLaMA2-7B-Chat w/ DeepEdit (Ours) 84.1
+Table 7: Specificity Score on the dataset CounterFact.
+Our DeepEdit outperforms all the baseline KE methods in terms of the Specificity Score, which validates that our DeepEdit causes less performance degradation for LLMs than the baseline methods when the edited facts do not change the ground-truth labels.
+Appendix C Iteration-wise Visualization of Our DeepEdit
+We present the iteration-wise outputs of our DeepEdit on an instance in Figure 8.
+Appendix D Case Study
 Figure 9 gives a qualitative comparison example between MeLLo and our DeepEdit on MQuaKE-3k.
@@
 ”, MeLLo is unable to generate the helpful sub-question to find the final correct answer.
-In contrast, our DeepEdit, based on the constraints of [Conciseness], [Coherence], and [Receptiveness], correctly find the relevant new knowledge “Bucharest is located in the continent of North America.” and obtain the correct answer North America with respect to the new knowledge.
+In contrast, our DeepEdit, effectively controls LLMs to correctly find the relevant new knowledge “Bucharest is located in the continent of North America.” and obtain the correct answer North America with respect to the new knowledge.
 Taking a closer look, we observe that our DeepEdit’s induced reasoning is much more succinct than MeLLo, since we effectively filter out the irrelevant and redundant information during decoding.
-Because longer and noisy textual information can lead to more serious hallucinations and higher difficulty of reasoning, our DeepEdit exhibits a stronger capability of KE for LLMs.
-Overall, our DeepEdit method enforces LLMs’ outputs to follow our decoding constraints so as to soundly incorporate new knowledge in reasoning, which leads to the correct predictions in Figure 8.
-6 Conclusion
-We explore a new paradigm for knowledge editing of black-box LLMs: decoding with constraints.
-We have designed a new decoding method DeepEdit to improve the knowledge editing of LLMs.
-We enforce the LLM’s outputs to satisfy the desired properties with decoding at the reasoning step level.
-Based on the DeepEdit, we can improve the LLMs’ reasoning with new knowledge though higher conciseness, coherence, pertinence, and receptiveness.
-Our DeepEdit controls LLMs’ reasoning to soundly incorporate new knowledge by satisfying our constraints.
-In addition to DeepEdit, we provide two new benchmarks: MQuAKE-2002 and MQuAKE-hard to provide more precise and challenging evaluation for KE methods respectively.
-Extensive experiments demonstrate the effectiveness of our DeepEdit.
-Future work includes analyzing the effects of other constraints that can improve knowledge editing for large language models.
+Because longer and noisy textual information can lead to more serious hallucinations and higher difficulty of reasoning, our DeepEdit exhibits a more effective KE than MeLLo with its succinct reasoning process controlled by constraints.
+Overall, our DeepEdit enhances LLMs to follow our decoding constraints so as to soundly incorporate new knowledge in multi-step reasoning, which leads to the correct predictions in Figure 9.

### v4

@@
 Abstract
-Answering multi-hop questions involving new knowledge is a challenging task in evaluating large language models’ (LLMs) knowledge editing (KE) methods.
-This task is rather difficult because the LLMs’ hallucinations on new knowledge would harm the logical coherence of LLMs’ multi-hop reasoning and lead to incorrect answers.
+How to edit the knowledge in multi-step reasoning has become the major challenge in the knowledge editing (KE) of large language models (LLMs).
+The difficulty arises because the hallucinations of LLMs during multi-step reasoning often lead to incorrect use of new knowledge and incorrect answers.
 To address this issue, we design decoding constraints to “regulate” LLMs’ reasoning, enhancing logical coherence when incorporating new knowledge.
-We incorporate the constraints into a new KE framework: DeepEdit ( Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs to generate coherent reasoning chains with new knowledge through a depth-first search.
+We propose a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs’s ability to generate coherent reasoning chains with new knowledge through depth-first search.
 Our search selects the most important knowledge that satisfies our constraints as the reasoning step to efficiently increase the reasoning depth.
@@
 1 Introduction
-Knowledge editing (KE) aims to enable large language models (LLMs) to produce outputs in accordance with the knowledge that is new or conflicted with the knowledge memorized by the LLMs.
-A challenging task to evaluate KE is the multi-hop question answering with new knowledge Zhong et al.
-(2023).
-In this task, LLMs should answer the questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
+Knowledge editing (KE) aims to enable large language models (LLMs) to incorporate new or updated knowledge, in contrast to relying solely on outdated parametric knowledge.
+Multi-hop question answering is the challenging task to evaluate KE Zhong et al.
+(2023), which requires models to answer questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
 Figure 1: (a) An instance of multi-hop question answering with new knowledge.
@@
 Figure 2: Our DeepEdit controls LLMs’ reasoning to follow the constraints of Conciseness, Coherence, Receptiveness, and Pertinence so as to soundly incorporate new knowledge into LLMs’ question answering.
-Recent literature has mainly sought to perform KE through in-context editing Zhong et al.
+Recent literature has mainly focused on performing KE through in-context editing Zhong et al.
 (2023); Cohen et al.
 (2023); Zheng et al.
-(2023), which puts new knowledge in the input context to fulfill KE.
-However, the input-conflicting hallucinations on the in-context new knowledge would harm the logical coherence of multi-step reasoning and lead to incorrect answersZhang et al.
+(2023), which puts new knowledge in the input context to fulfill KE (see Figure 1(b)).
+What new knowledge to use and how to utilize new knowledge is decided by the LLM generators’ implicit reasoning.
+However, the LLM generator’s hallucinations against the in-context new knowledge can harm the logical coherence of multi-step reasoning and result in incorrect answers Zhang et al.
 (2023).
-Taking a closer look at the ground-truth output in Figure 1(a), we notice that: the adjacent reasoning steps should be logically coherent, and new knowledge should participate in the reasoning chain at appropriate positions.
-This raises a natural question:
-Can we directly control the decoding of LLMs’ to soundly incorporate the new knowledge?
+To get the correct answers to multi-hop questions with new knowledge, we should ensure that the adjacent reasoning steps are logically coherent, and new knowledge is properly integrated into reasoning (see Fig.
+Figure 1(a).
+In this work, we explore using decoding constraints and search algorithms to explicitly control LLMs’ reasoning with new knowledge.
 The central idea of this paper is to view KE as a problem of constrained decoding.
-We design the decoding constraints, by following which LLMs can soundly incorporate new knowledge in LLMs reasoning.
-Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted memorized knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 2).
+We design the decoding constraints that facilitate LLMs to soundly incorporate new knowledge in LLMs reasoning.
+Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted parametric knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 2).
 A multi-step reasoning chain that leads to the correct answer should satisfy the above constraints.
 We incorporate the above constraints into a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing).
-When producing every reasoning step, DeepEdit searches over the parametric and new knowledge to find the ones that satisfy our constraints, termed as valid step candidates.
-Among the valid candidates, DeepEdit takes the most important one as the new reasoning step for further reasoning, which is the (new) knowledge that is most helpful to answer the target question.
-Our depth-first search strategy efficiently increases the reasoning depth and enhances LLMs to produce the coherent reasoning chains involving new knowledge.
-The advantages of our DeepEdit on KE are three-folds.
-First, it does not need to retrain the LLMs like prior model editing methods Zhu et al.
-(2020), which is complex and expensive.
-Our DeepEdit works during inference with negligible costs.
-Second, our DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions.
-Third, our DeepEdit enhances LLMs to soundly utilize new knowledge to produce coherent reasoning chains in accordance with new knowledge.
-To provide more precise and challenging assessments of KE methods, we build two new benchmarks MQuAKE-2002, and MQuAKE-hard, which resolve the knowledge-conflicting annotation mistakes in the popular KE benchmark MQuAKE-3k Zhong et al.
+When producing every reasoning step, DeepEdit searches through the parametric and new knowledge to find the knowledge items that satisfy our constraints, termed as valid step candidates.
+Among the valid candidates, DeepEdit prioritizes more important ones as the new reasoning step for further reasoning.
+We comprehensively evaluate the importance of different candidates based on their semantic relationships and the prior that the new knowledge is more informative to LLMs than the parametric knowledge.
+Note that the classical depth-first search has a backtracking mechanism that enables us to traverse all the reasoning chains of valid reasoning steps Tarjan (1972).
+However, such backtracking is undoubtedly time-consuming and unnecessary after an answer is found.
+Therefore, we design an early-stopping mechanism to stop the depth-first search when an answer is found by LLMs.
+Such an early stopping significantly improves the reasoning efficiency by reducing redundant iterations, as validated by the qualitative analysis (see Section 3.3) and empirical results (see Section 5.4).
+The advantages of DeepEdit are two-fold.
+First, DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions.
+Second, DeepEdit improves both the receptiveness of new knowledge and the coherence of multi-step reasoning by directly controlling the LLMs’ reasoning with decoding constraints.
+Unlike previous KE methods that edit either the model parameters Zhu et al.
+(2020) or the input prompts Cohen et al.
+(2023), we directly control the decoding of LLMs to incorporate the new knowledge into outputs, which opens up a new direction for KE in LLMs.
+To provide precise and challenging assessments of KE methods, we build two new benchmarks, MQuAKE-2002, and MQuAKE-hard, which resolve the knowledge-conflicting annotation mistakes in the popular KE benchmark MQuAKE-3k Zhong et al.
 (2023).
-We evaluate DeepEdit on the KE benchmarks MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al.
+We evaluate DeepEdit on MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al.
 (2023); Meng et al.
 (2022a).
-Qualitatively, DeepEdit enhances LLMs to produce more succinct and faithful reasoning with new knowledge, as shown in Figure 9.
+Qualitatively, DeepEdit enhances LLMs to produce succinct reasoning with new knowledge, as shown in Figure 9.
 Quantitatively, DeepEdit improves the KE for popular LLMs by a significant margin OpenAI (2022); Touvron et al.
@@
 (2021), as well as rapid adaptation facilitated by a compact auxiliary editing network Mitchell et al.
-(2022a); De Cao et al.
+(2022); De Cao et al.
 (2021); Mitchell et al.
@@
 In-context editing lacks direct control of the LLMs’ outputs and how the new knowledge influences the outputs is not traceable.
-Different from them, we propose the new decoding method DeepEdit for KE, which effectively augments LLMs on KE during the decoding stage.
+Different from the prior work, we propose a new KE method DeepEdit, which effectively augments LLMs on KE with constrained decoding.
 We illustrate a comparison between DeepEdit and existing KE methods in Table 1.
-Specifically, in our work, we consider a challenging setting of knowledge editing: the multi-hop question answering (QA) with knowledge editing Zhong et al.
-(2023).
-In-context learning methods combined with an optional retrieval module have become a popular approach in handling multi-hop QA tasks Press et al.
-(2022); Yao et al.
-(2022); Khattab et al.
-(2022); Mitchell et al.
-(2022b).
-We propose four types of decoding constraints to enhance question answering with KE.
-Our work improves multi-hop QA with KE by applying constraints to the reasoning of LLMs, which directly controls LLMs’ outputs to follow new knowledge.
-Figure 3: When producing every reasoning step, we retrieve the edited facts that are close to the generated step in the semantic space (left).
-Only the edited facts that satisfy all the constraints, i.e., valid edited facts, are considered as the step candidates (right).
-KE Method Editing Editing Editing Model Parameters Input Prompts Outputs MEND Mitchell et al. (2021) MEMIT Meng et al. (2022c) ICE Cohen et al. (2023) MeLLo Zhong et al. (2023) DeepEdit (Ours)
-Table 1: Comparison between KE methods.
+Our DeepEdit outperforms the prior KE methods by editing the LLMs’ outputs directly, which paves the way to a new direction of LLMs’ KE.
 Constrained Decoding
@@
 We design the semantic constraints to soundly incorporate new knowledge in LLMs’ reasoning, and verify the complex semantic constraints with LLMs.
+Figure 3: When producing every reasoning step, we retrieve the edited facts that are close to the generated step in the semantic space (left).
+Only the edited facts that satisfy all the constraints, i.e., valid edited facts, are considered as the step candidates (right).
+KE Method Editing Editing Editing Model Parameters Input Prompts Outputs MEND Mitchell et al. (2021) MEMIT Meng et al. (2022c) ICE Cohen et al. (2023) MeLLo Zhong et al. (2023) DeepEdit (Ours)
+Table 1: Comparison between KE methods.
 3 Methodology
-Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) of large language models (LLMs) Zhong et al.
-(2023).
-This task considers the questions that correspond to a chain of facts including new knowledge.
-It assumes a database that stores all new facts, which LLMs should adhere to, and posits that only a few facts can influence the ground-truth answer to each question.
+Multi-hop question answering with new knowledge assumes a database that stores all new facts, which LLMs should adhere to, and posits that only a few facts can influence the ground-truth answer to each question.
 In this section, we will introduce how we integrate the decoding constraints into the design of a novel KE method, DeepEdit (Depth-first Search based Decoding for Knowledge Editing), to enhance the KE of black-box LLMs.
 3.1 Constraint Designing for Knowledge Editing
-On a LLM’s multi-step reasoning chain to answer a multi-hop question involving new knowledge, we find the following constraints on reasoning steps can help LLMs to give the correct answer:
+On an LLM’s multi-step reasoning chain to answer a multi-hop question involving new knowledge, we find the following constraints on reasoning steps can help LLMs to give the correct answer:
   1. 1.
@@
 We visualize different constraints in Figure 2.
-Intuitively, the above constraints are easy to follow for LLMs when there is no new knowledge.
-However, the LLMs’ utilization of new knowledge is affected by the hallucination, which makes the above constraints easy to break when incorporating new knowledge.
-In this work, we apply the above constraints to incorporate the new knowledge into LLMs’ decoding.
+The above constraints are easy to follow for LLMs when there is no new knowledge.
+However, the utilization of in-context new knowledge is influenced by the LLMs’ hallucinations, which can easily break the aforementioned constraints.
+In this work, we apply the above constraints to incorporate the new knowledge into LLMs’ reasoning.
 3.2 Constraint Verification on Every Step
 We term the reasoning steps that satisfy all the constraints as valid steps.
-Because an invalid reasoning step could mislead LLMs into generating incorrect subsequent steps, we verify the constraints during the decoding of each reasoning step.
+Because an invalid reasoning step would mislead LLMs into generating incorrect subsequent steps, we verify the constraints at the decoding of each reasoning step.
 At every iteration, we take both the parametric and new knowledge as the step candidates.
-First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be useful to answer the target question.
+First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be most useful to answer the target question.
 Second, we retrieve $N$ new knowledge closest to the generated reasoning step.
-The distance between knowledge is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al.
+The distance between knowledge and the generated step is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al.
 (2019).
@@
 We take these $N+1$ knowledge as the reasoning step candidates, with examples shown in Figure 3.
-We verify the constraints on the above step candidates to find the valid step candidates.
+We verify the constraints on the aforementioned step candidates to identify the valid ones.
 In principle, every constraint’s verification can be seen as a binary classification problem on a pair of sentences.
@@
 In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers.
-Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from the instances of the MQuAKE dataset Zhong et al.
+Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from MQuAKE Zhong et al.
 (2023).
@@
 At every iteration, there may exist more than valid step candidates.
-We can take every one of them as the new reasoning step in parallel and make further reasoning based on each of them separately.
-This design can be seen as a breadth-first search of the LLMs’ multi-hop reasoning chain incorporating new knowledge.
+We can consider each valid step candidate simultaneously as the new reasoning step and make further reasoning based on each of them separately.
+This design can be seen as a breadth-first search of a valid multi-hop reasoning chain.
 It is undoubtedly time-consuming because the time complexity grows exponentially with the number of reasoning steps.
 This raises a natural question:
-Can we more efficiently expand the depth of reasoning chains during decoding?
+Can we increase the depth of reasoning chains more efficiently than the breadth-first search?
 Our answer is yes.
-We propose to take the most important step candidate as the reasoning step for further reasoning.
+We propose prioritizing more important step candidates as the reasoning step for further reasoning.
 Actually, different valid step candidates are not equally important.
-First, the edited facts should be more important than the generated step, since the former generally contains new knowledge unknown by LLMs.
-Second, the edited facts that are closer to the generated step are more important, which are more helpful in answering the target question due to their relevance to the LLM’s generated reasoning context.
+First, the edited facts should be more important than the generated step, since the former generally contains new knowledge unknown to LLMs.
+Secondly, the edited facts that are closer to the generated step are more important.
+They are more helpful in answering the target question due to their higher relevance to the context generated by the LLM’s reasoning.
 Based on the above prior, in every iteration, we select the most important knowledge as the next reasoning step and do further reasoning based on it, as shown in Figure 4.
@@
 Then we take the most important candidate as the next reasoning step to efficiently increase the reasoning depth.
+Figure 5: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k.
+In every iteration, the retrieved new knowledge (highlighted in green) is ranked from top to bottom as step candidates based on their semantic distances to the generated step, in ascending order.
 We name our method as DeepEdit: depth-first search-based constrained decoding for knowledge editing.
-DeepEdit takes the most important step candidate as the new reasoning step at every iteration and efficiently produces the multi-hop reasoning incorporating new knowledge.
-We summarize DeepEdit in Algorithm 1 and visualize it in Figure 1.
-Furthermore, we present the iteration-wise outputs of our DeepEdit in Figure 8.
+If we follow the classical depth-first search algorithm Tarjan (1972), we should include a backtracking mechanism to traverse all the reasoning chains of valid steps.
+However, such backtracking is undoubtedly time-consuming and unnecessary after an answer is found.
+Therefore, we design an early-stopping mechanism to stop the search when an answer is found by LLMs.
+We visualize DeepEdit in Figure 1 the example iteration-wise outputs in Figure 5.
 Assume the number of reasoning hops as $h$, and the number of valid step candidates per iteration as $c$.
-The time complexity of breadth-first search is $c^{h}$.
-In contrast, our DeepEdit has a much lower time complexity of $\mathcal{O}(h)$ thanks to our depth-first search design considering the importance of different candidates.
-Empirical results validate DeepEdit’s superior efficiency compared with the breadth-first search.
+The time complexity of breadth-first search is $\mathcal{O}(c^{h})$.
+In contrast, thanks to our depth-first search design with the early-stopping mechanism, our model has a significantly lower time complexity of $\mathcal{O}(h)$ in the best case.
+Empirical results validate DeepEdit’s superior efficiency compared with the breadth-first search (see Section 5.4).
 The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al.
 (2023).
-Different from them, our DeepEdit’s depth-first search is designed to efficiently find the multi-hop reasoning chain incorporating new knowledge.
-Our DeepEdit is free to be applied to all black-box LLMs: it bypasses the requirement of accessing the token-wise distributions because all the decoding actions are defined at the step level.
-Algorithm 1 DeepEdit: Depth-first Search based Constrained Decoding for Knowledge Editing.
-Input: The query question $q$. The prompt $p$. The set of new knowledge $\mathcal{K}$. The number of retrieved new knowledge $N$. Maximum number of reasoning depth $r$.
-Output: The answer $a$ to the query question.
-1:for $i\leftarrow 1$ to $r$ do
-2:     $o\leftarrow\mbox{LLM}(p)$
-3:     if $o$ is the answer then
-4:         return $o$
-5:     $\hat{\mathcal{K}}\leftarrow\mbox{search}(o,\mathcal{K},n)$triangleright retrieve the $n$ semantically similar new knowledge to $o$, ranked by the distance to $o$ in the ascending order.
-6:     for $j\leftarrow 1$ to $n$ do
-7:         $k\leftarrow\hat{\mathcal{K}}[i]$
-8:         $nk\leftarrow\mbox{False}$triangleright $nk$ is the boolean variable to decide whether some new knowledge is valid.
-9:         if ($k$ satisfies all the constraints) then
-10:              $nk\leftarrow\mbox{True}$
-11:              $p\leftarrow p+k$
-12:              break
-13:     if not $nk$ then
-14:         $p\leftarrow p+o$      return $o$
+Different from them, our DeepEdit’s depth-first search is designed to efficiently find the valid multi-hop reasoning chain incorporating new knowledge.
+Our DeepEdit is free to be applied to all black-box LLMs: it bypasses the requirement of accessing the token-wise distributions because our decoding strategy is defined at the reasoning step level.
 4 New Benchmarks for More Precise and Challenging Evaluation of KE
@@
 The best KE result on every LLM is highlighted in bold font.
-LLM as Constraint Verfier Conciseness Coherence Receptiveness Pertinence LLaMA2-7B-Chat 82.3 77.9 86.3 78.2 GPT-3.5-Turbo-Instruct 92.8 91.5 92.2 86.7
+LLM as Constraint Verfier CON COH REC PER
+LLaMA2-7B-Chat 82.3 77.9 86.3 78.2
+GPT-3.5-Turbo-Instruct 92.8 91.5 92.2 86.7
 Table 4: Experimental results (verification accuracy; %) of our verifiers for different constraints.
@@
 Figure 8: Test accuracy (%) of question answering with new knowledge with different KE batch sizes. We present the performance of MeLLo and our DeepEdit applied on GPT-3.5-Turbo-Instruct.
-Figure 8: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k.
-The chosen step candidate in every iteration that satisfies all the constraints are in green.
-The dropped step candidates are followed by the constraints that they do not satisfy.
 Figure 9: A case study for MeLLo and our DeepEdit on the dataset MQuAKE-3k. The new knowledge are in green. We report the prediction of Text-Davinci-003 with MeLLo and our DeepEdit.
@@
 Our DeepEdit outperforms all the baseline KE methods in terms of the Specificity Score, which validates that our DeepEdit causes less performance degradation for LLMs than the baseline methods when the edited facts do not change the ground-truth labels.
-Appendix C Iteration-wise Visualization of Our DeepEdit
-We present the iteration-wise outputs of our DeepEdit on an instance in Figure 8.
-Appendix D Case Study
+Appendix C Case Study
 Figure 9 gives a qualitative comparison example between MeLLo and our DeepEdit on MQuaKE-3k.
@@
 Overall, our DeepEdit enhances LLMs to follow our decoding constraints so as to soundly incorporate new knowledge in multi-step reasoning, which leads to the correct predictions in Figure 9.
+Prompt Text Generation Prompt of DeepEdit: Question: What is the capital of the country where Plainfield Town Hall is located?
+Thoughts with New Knowledge: Plainfield Town Hall is located in the country of the United States of America.# The capital of United States is Washington, D.C.# Washington, D.C.
+is the answer.
+Answer: Washington, D.C.
+Question: In which country is the company that created Nissan 200SX located?
+Thoughts with New Knowledge: Nissan 200SX was created by Nissan.# Nissan is located in the country of Japan.# Japan is the answer.
+Answer: Japan Question: Which continent is the country where the director of "My House Husband: Ikaw Na!" was educated located in?
+Thoughts with New Knowledge: The director of "My House Husband: Ikaw Na!" is Jose Javier Reyes.# Jose Javier Reyes was educated at De La Salle University.# De La Salle University is located in the country of Philippines.# Philippines is located in the continent if Asia.# Asia is the answer.
+Answer: Asia
+Figure 10: Text Generation Prompt of DeepEdit Zhong et al. (2023)
+Prompt Text Generation Prompt of MeLLo: Question: What is the capital city of the country of citizenship of Ivanka Trump’s spouse?
+Subquestion: Who is Ivanka Trump’s spouse?
+Generated answer: Ivanka Trump’s spouse is Jared Kushner.
+Retrieved fact: Ivanka Biden is married to Samantha Cameron.
+Retrieved fact does not contradict to generated answer, so the intermediate answer is: Jared Kushner Subquestion: What is the country of citizenship of Jared Kushner?
+Generated answer: The country of citizenship of Jared Kushner is United States.
+Retrieved fact: Jared Kushner is a citizen of Canada.
+Retrieved fact contradicts to generated answer, so the intermediate answer is: Canada Subquestion: What is the capital city of Canada?
+Generated answer: The capital city of Canada is Ottawa.
+Retrieved fact: The capital city of United States of America is Seattle.
+Retrieved fact does not contradict to generated answer, so the intermediate answer is: Ottawa Final answer: Ottawa Question: Who is the head of state of the country where Rainn Wilson holds a citizenship?
+Subquestion: What is the country of citizenship of Rainn Wilson?
+Generated answer: Rainn Wilson is a citizen of United States of America.
+Retrieved fact: Rainn Wilson is a citizen of Croatia.
+Retrieved fact contradicts to generated answer, so the intermediate answer is: Croatia Subquestion: What is the name of the current head of state in Croatia?
+Generated answer: The head of state of Croatia is President Zoran Milanović.
+Retrieved fact: The name of the current head of state in Croatia is Kolinda Grabar-Kitarović.
+Retrieved fact contradicts to generated answer, so the intermediate answer is: Kolinda Grabar-Kitarović Final answer: Kolinda Grabar-Kitarović Question: Who is the spouse of the head of state in United States of America?
+Subquestion: Who is the head of state in United States of America?
+Generated answer: The US president is Donald Trump.
+Retrieved fact: The head of state in United States of America is Joe Biden.
+Retrieved fact contradicts to generated answer, so the intermediate answer is: Joe Biden Subquestion: Who is the spouse of Joe Biden?
+Generated answer: The spouse of Joe Biden is Jill Biden.
+Retrieved fact: The spouse of Joe Bill is Evan Austin.
+Retrieved fact does not contradict to generated answer, so the intermediate answer is: Jill Biden Final answer: Jill Biden Question: On which continent is the country of citizenship of the founder of the manufacturer of iPhone 5 situated?
+Subquestion: Which company is iPhone 5 produced by?
+Generated answer: iPhone 5 is produced by Apple.
+Retrieved fact: The company that produced iPhone 5 is Iveco.
+Retrieved fact contradicts to generated answer, so the intermediate answer is: Iveco Subquestion: Who is the founder of Iveco?
+Generated answer: Iveco was founded by Giovanni Agnelli.
+Retrieved fact: House of Bonaparte was founded by Gustav I of Sweden.
+Retrieved fact does not contradict to generated answer, so the intermediate answer is: Giovanni Agnelli Subquestion: What is the country of citizenship of Giovanni Agnelli?
+Generated answer: Giovanni Agnelli is a citizen of Italy.
+Retrieved fact: Giovanni Agnelli is a citizen of Niger.
+Retrieved fact contradicts to generated answer, so the intermediate answer is: Niger.
+Subquestion: On which continent is Niger situated?
+Generated answer: Niger is situated on Africa.
+Retrieved fact: Kingdom of England is located in the continent of North America.
+Retrieved fact does not contradict to generated answer, so the intermediate answer is: Africa.
+Final answer: Africa
+Figure 11: Text Generation Prompt of MeLLo Zhong et al. (2023).

### v5

@@
 Abstract
-How to edit the knowledge in multi-step reasoning has become the major challenge in the knowledge editing (KE) of large language models (LLMs).
-The difficulty arises because the hallucinations of LLMs during multi-step reasoning often lead to incorrect use of new knowledge and incorrect answers.
-To address this issue, we design decoding constraints to “regulate” LLMs’ reasoning, enhancing logical coherence when incorporating new knowledge.
-We propose a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs’s ability to generate coherent reasoning chains with new knowledge through depth-first search.
-Our search selects the most important knowledge that satisfies our constraints as the reasoning step to efficiently increase the reasoning depth.
-In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which provide more precise and challenging assessments of KE approaches.
-Qualitatively, DeepEdit enables LLMs to produce succinct and coherent reasoning chains involving new knowledge.
+Incorporating new knowledge in multi-step reasoning has become a major problem in the research of knowledge editing (KE).
+LLMs often struggle to generate coherent and precise reasoning chains when incorporating new in-context knowledge, as they tend to deviate from relevant information and lose logical coherence during reasoning.
+To address this issue, it is crucial to apply direct controls over LLM outputs to ensure both logical coherence and effective integration of new knowledge into their reasoning processes.
+To this end, we propose DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing), a KE method designed to enhance LLMs’ reasoning capabilities with new knowledge.
+DeepEdit employs the decoding constraints that we carefully crafted to explicitly improve logical coherence and knowledge integration during reasoning.
+Additionally, we introduce a semantic-based importance estimator that ranks the relevance of reasoning steps to optimize the informativeness of output reasoning relative to the input question.
+Beyond DeepEdit, we propose two new KE benchmarks, MQuAKE-2002 and MQuAKE-hard, which offer more precise and challenging assessments of KE methodologies.
+Qualitative analysis shows that DeepEdit enables LLMs to generate succinct and coherent reasoning chains incorporating new knowledge.
 Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@
 1 Introduction
-Knowledge editing (KE) aims to enable large language models (LLMs) to incorporate new or updated knowledge, in contrast to relying solely on outdated parametric knowledge.
-Multi-hop question answering is the challenging task to evaluate KE Zhong et al.
+The research of knowledge editing (KE) aims to enable large language models (LLMs) to be such question-answering systems that the outputs are receptive to the new knowledge that is different from the LLMs’ parametric knowledge.
+Multi-hop question answering with new knowledge is currently the major task to evaluate KE Zhong et al.
 (2023), which requires models to answer questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
-Figure 1: (a) An instance of multi-hop question answering with new knowledge.
-(b) Prior KE methods, e.g., MeLLo Zhong et al.
-(2023), suffers from the hallucinations of LLM generators.
-(c) Our DeepEdit controls LLMs with decoding constraints to produce coherent reasoning chains involving new knowledge.
-Figure 2: Our DeepEdit controls LLMs’ reasoning to follow the constraints of Conciseness, Coherence, Receptiveness, and Pertinence so as to soundly incorporate new knowledge into LLMs’ question answering.
-Recent literature has mainly focused on performing KE through in-context editing Zhong et al.
+Figure 1: (a) An instance in KE evaluation benchmarks.
+(b,c) Prior KE methods that retrain LLMs with new knowledge (model editing) and adds new knowledge in the input prompt (in-context editing).
+(c) Our DeepEdit directly controls LLMs’ outputs to select the valid and most important (new) knowledge for reasoning.
+Existing literature has mainly focused on performing KE through model editing Mitchell et al.
+(2021); Meng et al.
+(2022b) and in-context editing Zhong et al.
 (2023); Cohen et al.
 (2023); Zheng et al.
-(2023), which puts new knowledge in the input context to fulfill KE (see Figure 1(b)).
-What new knowledge to use and how to utilize new knowledge is decided by the LLM generators’ implicit reasoning.
-However, the LLM generator’s hallucinations against the in-context new knowledge can harm the logical coherence of multi-step reasoning and result in incorrect answers Zhang et al.
 (2023).
-To get the correct answers to multi-hop questions with new knowledge, we should ensure that the adjacent reasoning steps are logically coherent, and new knowledge is properly integrated into reasoning (see Fig.
+The former uses retraining to inject new knowledge into LLMs’ weights while the latter adds new knowledge to the input prompt to fulfill KE (see Figure 1(b)).
+Both of them lack direct control of LLMs’ outputs, which is essential to enhance LLMs’ multi-step reasoning incorporating new knowledge (see Figure 2).
+A satisfactory multi-step reasoning chain with new knowledge should consist of coherent steps and place new knowledge at appropriate positions (see Fig.
 Figure 1(a).
-In this work, we explore using decoding constraints and search algorithms to explicitly control LLMs’ reasoning with new knowledge.
-The central idea of this paper is to view KE as a problem of constrained decoding.
-We design the decoding constraints that facilitate LLMs to soundly incorporate new knowledge in LLMs reasoning.
+In this work, we explore to propose a new constrained decoding algorithm that enhances LLMs’ to produce such reasoning chains.
+First, we design the decoding constraints that facilitate LLMs to soundly incorporate new knowledge in multi-step reasoning.
 Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted parametric knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 3).
-A multi-step reasoning chain that leads to the correct answer should satisfy the above constraints.
-We incorporate the above constraints into a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing).
+Reasoning without New Knowledge Question: The writer of the novel “1984” died in which continent?
+Reasoning Steps: The writer of the novel "1984" is George Orwell.
+George Orwell died in the city of London.
+London is in the continent of Europe.
+Answer: Europe ✓
+Reasoning with New Knowledge in Context Question: Assume George Orwell died in the city of Bucharest; The writer of “Harry Potter” is Stephen King; Bucharest is located in the continent of North America; Berlin is located in the continent of Asia.
+The writer of the novel “1984” died in which continent?
+Reasoning Steps: The writer of the novel "1984" is George Orwell.
+The writer of “Harry Potter” is Stephen King.
+George Orwell died in the city of Bucharest.
+Berlin is located in the continent of Asia.
+Answer: Asia ✗
+Reasoning with New Knowledge in Decoding (Ours) Question: The writer of the novel “1984” died in which continent?
+Reasoning Steps: The writer of the novel "1984" is George Orwell.
+George Orwell died in the city of Bucharest.
+Bucharest is located in the continent of North America.
+Answer: North America ✓
+Figure 2: (Upper) It is easy for LLMs to generate coherent reasoning chains without new knowledge.
+(Middle) Given new knowledge in context, LLMs can hardly generate coherent reasoning chains due to hallucinations on new knowledge.
+As a result, both relevant and irrelevant knowledge participate in the reasoning.
+(Lower) We do not put new knowledge in context but let them directly contribute to the decoding, which leads to coherent and precise reasoning chains.
+We incorporate the above constraints into a new KE method: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing).
 When producing every reasoning step, DeepEdit searches through the parametric and new knowledge to find the knowledge items that satisfy our constraints, termed as valid step candidates.
@@
 We comprehensively evaluate the importance of different candidates based on their semantic relationships and the prior that the new knowledge is more informative to LLMs than the parametric knowledge.
-Note that the classical depth-first search has a backtracking mechanism that enables us to traverse all the reasoning chains of valid reasoning steps Tarjan (1972).
-However, such backtracking is undoubtedly time-consuming and unnecessary after an answer is found.
+Classical depth-first search has a backtracking mechanism that enables us to traverse all the reasoning chains of valid reasoning steps Tarjan (1972).
+Such backtracking is time-consuming and unnecessary after an answer is found.
 Therefore, we design an early-stopping mechanism to stop the depth-first search when an answer is found by LLMs.
 Such an early stopping significantly improves the reasoning efficiency by reducing redundant iterations, as validated by the qualitative analysis (see Section 3.3) and empirical results (see Section 5.4).
-The advantages of DeepEdit are two-fold.
-First, DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions.
-Second, DeepEdit improves both the receptiveness of new knowledge and the coherence of multi-step reasoning by directly controlling the LLMs’ reasoning with decoding constraints.
+DeepEdit is flexibly applicable to any black-box LLM that only requires access to output texts.
+Besides, DeepEdit improves both the receptiveness of new knowledge and the coherence of multi-step reasoning by directly controlling the LLMs’ reasoning with decoding constraints.
 Unlike previous KE methods that edit either the model parameters Zhu et al.
 (2020) or the input prompts Cohen et al.
-(2023), we directly control the decoding of LLMs to incorporate the new knowledge into outputs, which opens up a new direction for KE in LLMs.
+(2023), we directly control the output of LLMs to incorporate the new knowledge, which opens up a new direction for KE in LLMs.
 To provide precise and challenging assessments of KE methods, we build two new benchmarks, MQuAKE-2002, and MQuAKE-hard, which resolve the knowledge-conflicting annotation mistakes in the popular KE benchmark MQuAKE-3k Zhong et al.
@@
 (2023).
+Figure 3: Our DeepEdit controls LLMs’ reasoning to follow the constraints of Conciseness, Coherence, Receptiveness, and Pertinence so as to soundly incorporate new knowledge into LLMs’ question answering.
 2 Related Work
@@
 Different from the prior work, we propose a new KE method DeepEdit, which effectively augments LLMs on KE with constrained decoding.
-We illustrate a comparison between DeepEdit and existing KE methods in Table 1.
 Our DeepEdit outperforms the prior KE methods by editing the LLMs’ outputs directly, which paves the way to a new direction of LLMs’ KE.
@@
 We design the semantic constraints to soundly incorporate new knowledge in LLMs’ reasoning, and verify the complex semantic constraints with LLMs.
-Figure 3: When producing every reasoning step, we retrieve the edited facts that are close to the generated step in the semantic space (left).
-Only the edited facts that satisfy all the constraints, i.e., valid edited facts, are considered as the step candidates (right).
-KE Method Editing Editing Editing Model Parameters Input Prompts Outputs MEND Mitchell et al. (2021) MEMIT Meng et al. (2022c) ICE Cohen et al. (2023) MeLLo Zhong et al. (2023) DeepEdit (Ours)
-Table 1: Comparison between KE methods.
 3 Methodology
@@
   1. 1.
-    Conciseness: Every step is identical.
+    Conciseness: Every step is unique.
   2. 2.
@@
 The closer the context of the new knowledge is to the generated step, the more likely it is that the new knowledge will be helpful in answering the target question.
-We take these $N+1$ knowledge as the reasoning step candidates, with examples shown in Figure 3.
+We take these $N+1$ knowledge as the reasoning step candidates.
 We verify the constraints on the aforementioned step candidates to identify the valid ones.
@@
 In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers.
-Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from MQuAKE Zhong et al.
+Every verifier’s prompt includes $D$ positive and $N$ negative demonstrations randomly sampled from MQuAKE Zhong et al.
 (2023).
@@
 3.3 Efficient Reasoning with Depth-first Search
-At every iteration, there may exist more than valid step candidates.
+At every iteration, there may exist more than one valid step candidates.
 We can consider each valid step candidate simultaneously as the new reasoning step and make further reasoning based on each of them separately.
@@
 (2023), a recent KE dataset that is designed to evaluate the KE methods on helping LLMs to answer multi-hop questions given new knowledge.
-Every instance in MQuAKE includes a multi-hop question and several edited facts, every of which can alter the ground-truth answer to the question.
+Every instance in MQuAKE includes a multi-hop question and several edited facts, each of which can alter the ground-truth answer to the question.
 4.1 Annotation Mistakes in One-Third Instances of MQuAKE
@@
 (2023) to set the KE batch size as 100 and full batch for evaluation.
-The KE batch size means the number of instances that provide new knowledge for retrieval when answering every question.
+The KE batch size means the number of instances that provide new knowledge for retrieval when answering each question.
 Compared methods.
@@
 Future work includes exploring other constraints that can improve knowledge editing for large language models.
-Acknowledgements
-This work was partially supported by a DARPA ANSR program FA8750-23-2-0004.
-The views and conclusions are those of the authors and should not reflect the official policy or position of DARPA or the U.S.
-Government.
+Limitations
+DeepEdit is an early effort at conducting knowledge editing in the decoding stage.
+There still exist spaces for improving DeepEdit.
+For example, can we have more decoding constraints to improve knowledge editing?
+And can we develop more advanced decoding strategies to make LLMs follow the constraints?
+We hope the future work can answer above questions and further improve knowledge editing to serve more real-world applications.
+Ethical Considerations
+Ethical considerations are of utmost importance in our research endeavors.
+In this paper, we strictly adhere to ethical principles by exclusively utilizing open-source datasets and employing various models that are either open-source or widely recognized in the scientific community.
+We are committed to upholding ethical standards throughout the research process, prioritizing transparency, and promoting the responsible use of technology for the betterment of society.
 Appendix A Analysis on Knowledge Editing Batch Size
@@
 Overall, our DeepEdit enhances LLMs to follow our decoding constraints so as to soundly incorporate new knowledge in multi-step reasoning, which leads to the correct predictions in Figure 9.
+Appendix D Discussion with MeLLo
+MeLLo Zhong et al.
+(2023) utilizes a prompt engineering based greedy search to produce a reasoning chain that is receptive to new knowledge.
+At every iteration, MeLLo retrieves a new knowledge item that is semantically close to the generated reasoning step.
+It assesses whether the retrieved new knowledge is conflicted with the generated step.
+If so, it takes the new knowledge as next reasoning step, or the generated step otherwise.
+Below, we compare the differences between MeLLo and our DeepEdit from different perspectives:
+  • Generation Prompt.
+As shown in Fig.
+Figure 10 and Figure 11, MeLLo’s generation prompt is much more complex than our DeepEdit.
+MeLLo leads to much longer context than our DeepEdit for every inference, which induces more serious hallucinations and time/computational costs.
+  • Greedy Search versus Depth-First Search.
+MeLLo employs the greedy search to find the reasoning chain consisting of both parametric and new knowledge.
+If there is error on any reasoning step, the final answer cannot be gotten without the backtracking mechanism.
+We employ the depth-first search to look for satisfactory reasoning chains with new knowledge, of which the backtracking mechanism provides a fault-tolerance mechanism to complement the mistakes happening in any reasoning steps.
+  • New Knowledge Selection How and where to use the new knowledge in the reasoning chain is the key to getting the correct answer with new knowledge.
+MeLLo only uses a conflict assessment defined in the prompt template to determine whether to use a new knowledge item.
+In contrast, our DeepEdit comprehensively evaluates different new knowledge using explicit constraints and a semantic-based importance estimator.
 Prompt Text Generation Prompt of DeepEdit: Question: What is the capital of the country where Plainfield Town Hall is located?
@@
 Answer: Asia
-Figure 10: Text Generation Prompt of DeepEdit Zhong et al. (2023)
+Figure 10: Text Generation Prompt of DeepEdit (Ours).
 Prompt Text Generation Prompt of MeLLo: Question: What is the capital city of the country of citizenship of Ivanka Trump’s spouse?